### v1

Abstract
Scaling up language models to billions of parameters has opened up possibilities for in-context learning, allowing instruction tuning and few-shot learning on tasks that the model was not specifically trained for. This has achieved breakthrough performance on language tasks such as translation, summarization, and question-answering. Furthermore, in addition to these associative “System 1” tasks, recent advances in Chain-of-thought prompt learning have demonstrated strong “System 2” reasoning abilities, answering a question in the field of artificial general intelligence whether LLMs can reason.
The field started with the question whether LLMs can solve grade school math word problems. This paper reviews the rapidly expanding field of prompt-based reasoning with LLMs. Our taxonomy identifies different ways to generate, evaluate, and control multi-step reasoning. We provide an in-depth coverage of core approaches and open problems, and we propose a research agenda for the near future. Finally, we highlight the relation between reasoning and prompt-based learning, and we discuss the relation between reasoning, sequential decision processes, and reinforcement learning. We find that self-improvement, self-reflection, and some metacognitive abilities of the reasoning processes are possible through the judicious use of prompts. True self-improvement and self-reasoning, to go from reasoning with LLMs to reasoning by LLMs, remains future work.
1 Introduction
Transformer-based Large Language Models (LLMs) that are trained on large datasets have achieved breakthrough performance at next token prediction (Vaswani et al., 2017; Radford et al., 2019; Wei et al., 2022a); they are very good at natural language understanding (GLUE, SQUAD, Xsum) (Wang et al., 2018, 2019; Rajpurkar et al., 2016; Narayan et al., 2018), translation (Kocmi et al., 2022; Papineni et al., 2002; Sennrich et al., 2015), question answering (Tan et al., 2023), and other System 1 tasks (Kahneman, 2011).111In his book Thinking, fast and slow, a bestseller on human psychology, Daniel Kahneman described System 1 thinking as a near-instantaneous process; it happens automatically, intuitively, and with little effort. It is driven by instinct and experiences. System 2 thinking is slower and requires more effort. It is conscious and logical. The automatic operations of System 1 generate surprisingly complex patterns of ideas, but only the slower System 2 can construct thoughts in an orderly series of steps. In the LLM literature the terms are often used as shorthand to distinguish single-step associative tasks, from multi-step reasoning tasks, despite the fact that language tasks such as question answering and translation may require some “slow” thinking. The success of ChatGPT (Ouyang et al., 2022) has taken the world by storm.
Transformer-based generative language models whose size is beyond hundreds of billions parameters are not only very good at language generation, they also enable new type of machine learning, called in-context learning (Brown et al., 2020). In-context learning, also known as prompt-based learning, occurs only in LLMs beyond a certain size (hundreds of billions of parameters) that are sufficiently rich (Wei et al., 2022a). In-context learning is inference time, prompt-based, few-shot learning, where model parameters are not trained or fine-tuned.
System 1 tasks, such as associative language tasks, are easily solved by LLMs with prompt-based learning, as the many school children around the world that use ChatGPT daily can attest. (Although the problems are too often not solved correctly, just fluently, when the model’s association powers lead to hallucination (Huang et al., 2023).) On the other hand, System 2 tasks, such as grade school math word problems, are more difficult for LLMs(Cobbe et al., 2021). To solve math word problems we need to break down the problem in multiple reasoning steps. Spurred-on by the impressive performance on System 1 tasks, much research has focused on understanding the reason for the poor performance of LLMs on System 2 tasks, and how it can be improved.
Among this research, the Chain-of-thought experiment (Wei et al., 2022b) stands out. This work, and subsequently Kojima et al. (2022), showed that adding a simple instruction to the prompts, Let’s think step by step, can provoke an LLM to perform the required intermediate reasoning steps, achieving a surprising jump in performance. The Chain-of-thought paper is a breakthrough in the field of reasoning with LLMs. Much exciting work has been published that builds on this work.
Grade school math word problems started the research into LLM-reasoning, with the GSM8K benchmark (Cobbe et al., 2021). In our survey we discuss papers based on this benchmark, and directly-related follow up work on reasoning. We focus on prompt-based approaches. We survey the recent literature using a straightforward taxonomy.
Although the field has only recently started, the jump in performance on reasoning has excited artificial intelligence and society alike. We provide a research agenda with opportunities for future research. At the end of this survey, we also discuss connections to other fields, such as self-reflection, metacognition (or thinking about thinking, see for example Dunlosky and Metcalfe (2008)), and the motivation towards artificial general intelligence.
Our contributions are:
A survey of relevant approaches in prompt-based reasoning (grade school math word problems and closely related domains) in large language models, including a research agenda.
A taxonomy based on regular reasoning literature (step generation, step evaluation, and control of reasoning steps).
This survey is organized as follows. Section 2 summarizes the most relevant developments in LLMs, including in-context learning. Of great importance are the benchmarks that are used in this field. We discuss these in Section 3, followed by our method for scoping and selecting of papers in Section 4. Next, in Section 5 we provide a taxonomy of the field, where we discuss the approaches in detail. Then, in Section 6 we discuss our findings in a broader perspective. We also discuss the relation between reasoning and work on self-reflection and metacognition. This section concludes with an agenda for future research. Finally, Section 7 concludes the survey.
2 Background: Reasoning with LLMs
Before we dive into the works on reasoning, we review some background terminology on LLMs. Our overview is brief. Excellent surveys on LLMs are, for example, Minaee et al. (2024) and Zhao et al. (2023). We discuss the generic training pipeline for LLMs, we discuss how in-context learning works, and we discuss the reasoning pipeline. We start with the generic language model training pipeline.
2.1 Training Pipeline Language Model
LLMs are typically constructed in a sequence of stages, from data preparation, through training, to inference. The training pipeline for most LLMs is quite elaborate. We will now list a pipeline of the most common stages, based on the survey by Minaee et al. (2024).
1.
Acquire a large, general, unlabeled, high-quality text corpus. Some considerations on the selection of the texts are discussed in Brown et al. (2020).
2.
Pretrain the transformer model (Vaswani et al., 2017) on this large corpus. This step yields a generalist model. The pretraining is done using a self-supervised approach on the unlabeled dataset (text corpus).
3.
Finetune the general model to a specific (narrow) task. This can be done using supervised-learning with a new labeled dataset consisting of prompts and answers (supervised finetuning, SFT) (Wei et al., 2022a; Minaee et al., 2024), specific for the task at hand. (A small number of papers in this survey work in the finetuning stage.)
4.
Instruction tuning is a form of finetuning on a labeled dataset of instruction prompts and corresponding outputs, to improve instruction following, and thus the usefulness of models.
5.
Align the finetuned model with user expectations (preference alignment). The goal of this stage is to improve the model to give more ethically and socially acceptable answers. The machine learning method that is used in this stage can be, for example, Reinforcement Learning with Human Feedback (Ouyang et al., 2022) or Direct Preference Optimization (Rafailov et al., 2024).
- 6.
7.
Inference & In-context learning can be used to train the model to provide the correct answers without changing parameters (Dong et al., 2022; Brown et al., 2020). By providing a prompt that contains a small number of examples together with a question, prompt learning is a form of few-shot learning. This is the stage in which most of the papers of this survey work, and that is familiar to all general users of ChatGPT.
Most of the reasoning methods that we discuss in this survey work in stage 7: in-context learning, using prompts for the LLM to perform a complex multi-step reasoning task. The following section provides a brief introduction to in-context learning.
2.2 In-Context Learning
In LLMs beyond hundreds of billions of parameters a new kind of learning has emerged, that is called in-context learning or prompt-learning (Brown et al., 2020). It occurs at inference time, and is often able to give good results with few examples; it is a form of few-shot learning. The large size of the model, containing rich and general knowledge is enabling this new type of few-shot learning (see Dong et al. (2022) for a survey).
A prompt, consisting of a piece of demonstration context, is concatenated with a query question, and is given to the language model for prediction (Liu et al., 2023). For example, when the task is emotion recognition in a social media post, “I missed the bus today,” can be followed by “I felt so [___]”. Alternatively, for translation, we could follow “I missed the bus today,” by “French: [___]” (Liu et al., 2023). The prompt contains background information that is recognized by the model, selecting the desired model context. In-context learning works when language models contain enough knowledge, allowing them to generalize on the examples provided in the prompt.
Prompts that contain a few examples are said to perform few-shot learning. Prompts that contain only instructions without examples are said to perform zero-shot learning.
In-context learning takes place at inference time, after the computationally intensive stages where parameters have been pretrained and finetuned, when the model is queried by the user to provide answers. No parameters are changed anymore with in-context learning. This is quite different from the common approach in supervised deep learning—or self-supervised deep learning—where large datasets are used during training to update model parameters with backward propagation in lengthy and costly training epochs (Goodfellow et al., 2016). Common approaches to few-shot learning, such as metalearning, do include training and finetuning of parameters to achieve generalization, and are computationally expensive (see, for example, Finn et al. (2017) or Huisman et al. (2021); Hospedales et al. (2021) for a survey).
Prompts provide a user-friendly interface to LLMs. The success of in-context learning tends to be quite sensitive to the way a prompt is formulated; a new field called prompt engineering has emerged to optimize the usefulness of in-context learning by learning how to make them do what we want (Radford et al., 2019; Wei et al., 2022a; Giray, 2023; Sahoo et al., 2024).
2.3 Reasoning Pipeline
Reasoning problems are also solved with a pipeline of stages. A typical approach to solving a complex problem is to subdivide it into smaller steps and solve those. This approach is related to divide and conquer (Bellman, 1966). New steps are (1) generated, (2) evaluated, and the number of steps that are generated and searched is (3) controlled in some way. The in-context reasoning approaches that we survey follow a general three-stage pipeline (Madaan et al., 2023):
1.
Generate: generation of steps by the model,
2.
Evaluate: evaluation of the predicted steps by an evaluator,
3.
Control: control of the number of steps that are generated and how deep ahead the reasoning process will look.
This three-stage pipeline will be the basis of our taxonomy. But first, we will look at benchmarks.
3 Benchmarks
Progress in artificial intelligence is measured by benchmarks. Benchmarks define the goal that researchers aim to achieve in their experiments. In natural language processing, a wide array of benchmarks exists to measure progress, such as on question answering (for example, CommonsenseQA (Talmor et al., 2018)), word prediction (for example, LAMBADA (Paperno et al., 2016)), translation (for example, WMT’22 (Kocmi et al., 2022)), language understanding (for example, GLUE (Wang et al., 2018, 2019)), and text summarization (for example, Xsum (Narayan et al., 2018)). Transformer architectures were first popularized by encoder models such as BERT (Devlin et al., 2018), for named entity recognition and classification tasks. Subsequently, decoder models such as GPT 2-4 (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023) showed impressive progress on natural language benchmarks.
The field of LLMs is quite active. Many different benchmarks exist, and listing a comprehensive overview of all relevant benchmarks is beyond the scope of this survey. We will mention relevant benchmarks for testing the reasoning abilities of LLMs. Following Wei et al. (2022b), these are all math word problem benchmarks. The benchmark that is most frequently associated with reasoning by LLMs is a dataset of grade school math word problems GSM8K (Cobbe et al., 2021). GSM8K was created by humans, with an aim of high quality, high diversity, moderate difficulty, and solutions in natural language. Other benchmarks are the SVAMP varying structures benchmarks (Patel et al., 2021), the ASDiv dataset of diverse math problems (Miao et al., 2021), the AQuA dataset of algebraic word problems (Ling et al., 2017), and the MAWPS benchmark (Koncel-Kedziorski et al., 2016).
We will now briefly discuss these benchmarks; the baseline performance that we quote is from Wei et al. (2022b).
GSM8K
To test reasoning skills, the Grade School Math problem dataset (GSM8K) was developed for testing LLMs (Cobbe et al., 2021). It consists of 8500 human-written math problems. Language models struggled to achieve good performance on this dataset (pre Chain-of-thought). An example of a math word task is:
Problem: Beth bakes 4, two dozen batches of cookies in a week. If these cookies are shared amongst 16 people equally, how many cookies does each person consume?
Answer: .
The baseline performance of GPT-3 175B is 15.6% accuracy. In comparison, the performance of Chain-of-thought is 46.9% accuracy.
ASDiv
The Academia Sinica Diverse MWP Dataset (ASDiv) (Miao et al., 2021) is specifically designed for high diversity in problem types, formats and difficulty levels. It consists of 2305 problems. An example problem is:
Problem: A sandwich is priced at . A cup of pudding is priced at . Tim bought 2 sandwiches and 4 cups of pudding. How much money should Tim pay?
Answer: .
The baseline performance of GPT-3 175B is 70.3% accuracy. The performance of Chain-of-thought is 71.3% accuracy.
MAWPS
The Math Word Problem Repository (MAWPS) (Koncel-Kedziorski et al., 2016) allows for the construction of datasets with particular characteristics by selecting different categories of problems. The dataset consists of 3320 problems. An example is:
Problem: Rachel bought two coloring books. One had 23 pictures and the other had 32. After one week she had colored 44 of the pictures. How many pictures does she still have to color?
Answer: .
The baseline performance of GPT-3 175B is 72.7% accuracy. The performance of Chain-of-thought is 87.1% accuracy.
SVAMP
The Simple Variations on Arithmetic Math word Problems dataset (SVAMP) was designed by Patel et al. (2021). It consists of 1000 problems, curated from variations of ASDiv-a (Miao et al., 2021) and MAWPS (Koncel-Kedziorski et al., 2016). An example problem is:
Problem: Jack had 8 pens and Mary had 5 pens. Jack gave 3 pens to Mary. How many pens does Jack have now?
Answer: .
The baseline performance of GPT-3 175B is 65.7% accuracy. In comparison, the performance of Chain-of-thought is 68.9% accuracy.
AQuA
The Algebraic Question Answering dataset (Ling et al., 2017) is a large dataset of 100,949 questions, answers, and rationales. The dataset is based on a combination of a smaller seed dataset and crowdsourcing. An example question is:
Question: Two trains running in opposite directions cross a man standing on the platform in 27 seconds and 17 seconds respectively and they cross each other in 23 seconds. The ratio of their speeds is: Options: A) 3/7 B) 3/2 C) 3/88 D) 3/8 E) 2/2
Answer: B.
The baseline performance of GPT-3 175B is 24.8% accuracy. The performance of Chain-of-thought is 35.8% accuracy.
There is a wide variety of benchmarks, and there is a wide variety of performance in benchmarks. Some are easily solvable by current LLMs, and some are significantly harder. Benchmark design is an important part of the field of reasoning in LLMs. Currently the GSM8K benchmark is popular; baseline model performance is weak, and reasoning prompts can substantially improve performance. As performance on GSM8K improves, different (harder) benchmarks will become popular.
4 Selection of Papers
The papers in this survey were selected as follows. Baseline LLMs have difficulty solving math word problems, specifically on benchmarks listed in the previous section. We take the ability to solve those benchmarks as a proxy for reasoning ability. We initially performed a literature search for papers that use these benchmarks, and that contain the search terms reasoning and large language model in their title or abstract. We also searched for papers that referenced the Chain-of-thought paper. The resulting papers were curated based on recency, relevance, substance, and novelty.
We favor recent papers (two years prior to the writing of the survey), related to the Chain-of-thought approach of generating intermediate reasoning steps, that solve tasks such as math word problems, and that work by prompt-based in-context learning. We also include some papers that work by finetuning or supervised learning that relate to, or inspire, the Chain-of-thought approaches. Furthermore, we include approaches outside math word problems that showed interesting approaches to reasoning, such as applications in coding and autonomous agents, because of their approach to grounding.
5 Prompt Generation, Evaluation and Control
This survey examines how an architecture that is good at System 1 tasks can be prompted to solve System 2 tasks. The Chain-of-thought paper showed how a simple command could prompt an LLM to perform reasoning steps, yielding much better performance in math word problems. Since then much research has further explored this approach, trying to build the ultimate general problem solver for System 1 and System 2 problems.
Following the pipeline of Section 2.3, the prompts must (1) generate the reasoning steps, (2) evaluate the answer to the steps, and (3) control the number of steps that are generated, the shape (or complexity) of the reasoning process must be controlled. We will now briefly discuss the three stages. Please refer to Figure 1 for a diagram of the different approaches for the generation, evaluation, and control of reasoning steps, and to Table 1.222We show the approaches in the Figure in their main category only. Some approaches show innovations in two categories, and are shown twice. (Since all approaches have a generation, an evaluation, and a control aspect, all could in principle occur three times—all three columns can be found in Table 1).
Prompt for Step Generation
The first order of business is to create a prompt that instructs the LLM to generate reasoning steps. The problem must be split into substeps. This can be achieved with a problem-specific prompt that contains elements of the problem, such as: “First calculate how many marbles Mary had originally, then how many her friend had, and finally how many they had together.”
In general, it is possible to prompt an LLM to fill in the blanks in a step-by-step fashion. In the papers that we will discuss, there are three main approaches for generating the step-by-step prompt. The prompt may be (1) handcrafted for the problem by the researchers (hand-written prompt), or (2) the prompt or prompts may come from an source that is external to the model, such as another model or a dataset (prompt using external knowledge), or (3) the model itself can be prompted to generate a (series of) prompt(s) to analyze the problem (model-generated prompt). As we will see, all three approaches have their advantages and disadvantages.
Generating the subproblem-steps is the first stage that is necessary for in-context learning to perform reasoning. Each paper in our survey performs at least this stage of the reasoning pipeline. In some of the early papers (around 2022) it is the only stage of the pipeline that is performed.
Prompt for Result Evaluation
After the prompt has been generated and the model has answered it, the next step in the reasoning pipeline is to evaluate the answer. Again, we see three main approaches for substep evaluation. First, the steps may be evaluated by (1) the model itself (self-assessment). Second, (2) an external program can be used to evaluate the steps. For example, when the steps are expressed as computer code, an external interpreter or compiler can be used to check the validity and the outcome (tool-based evaluation). Finally, (3) an external model can be used, LLM or otherwise. For example, in robotics, an external physics model can determine if certain actions are physically possible (external model validation).
Perform Control of Reasoning Steps
A reasoning process that consists of multiple steps is a sequential decision process (Littman, 1996). When a single chain of reasoning steps is generated, the control flow of the reasoning process is simple: greedily evaluate the first step and then the next one, if present. The control flow of the reasoning process may also be more intricate. Some reasoning problems can be divided into multiple subproblems. To execute, evaluate and combine the results of all substeps, a separate controller may be needed. This controller can be a prompt or an external algorithm.
Again, we distinguish three approaches. Most papers use (1) a greedy selection approach: a single prompt with a single chain of steps is generated, and these steps are directly executed and followed. The second approach (2) is to generate an ensemble strategy of reasoning steps, evaluate them, combine the individual results, and present them as the result of the ensemble. Finally, (3) a full tree-search or a reinforcement learning (RL) algorithm can be used as scaffolding. In this case, when a step is followed and evaluated, the LLM can roll back and try a different reasoning step. This is a breadth-first search approach (Plaat, 2020). Going further, a full reinforcement learning approach can be used (Sutton and Barto, 2018; Plaat, 2022) to find an optimal policy for the sequential decision process. A full Markov Decision Process of state, action, transition, and reward function is specified, and step control can become a process where prompts are generated dynamically.
| Approach | Domain | Step generation | Step evaluation | Step control |
| Scratchpad (Nye et al., 2021) | math word | hand-wr/supervised | - | greedy/1prompt |
| Chain-of-thought (Wei et al., 2022b) | math word | hand-written | - | greedy/1prompt |
| ZS-CoT (Kojima et al., 2022) | math word | hand-written | - | greedy/1prompt |
| Auto-CoT (Zhang et al., 2022) | math word | model-generated | - | clustering |
| Complexity (Fu et al., 2022) | math word | hand-written | self-consistency | greedy/1prompt |
| Self-ask (Press et al., 2022) | math word | external knowledge | LLM | multi-hop questions |
| Self-verification (Weng et al., 2022) | math word | hand-written | back-verify | ensemble |
| Self-consistency (Wang et al., 2022b) | math word | hand-written | majority | ensemble |
| Codex (Chen et al., 2021) | code | - | tool-based | - |
| Self-debugging (Chen et al., 2023) | code | hand-written | tool-based | greedy |
| Fun-search (Romera-Paredes et al., 2024) | code | hand-written | tool-based | evolutionary algorithm |
| LLaMEa (van Stein and Bäck, 2024) | code | hand-written | tool-based | evolutionary algorithm |
| MathPrompter (Imani et al., 2023) | math | hand-written | tool-based | ensemble |
| Program-of-thoughts (Chen et al., 2022) | math word | hand-written, Codex | Python+Consist. | decouple reason/compute |
| Program-aided-language (Gao et al., 2023) | math word | hand-written, Codex | NLP/Python | ensemble |
| Refiner (Paul et al., 2023) | math word | finetune | critic model | gen/crit feedback |
| Self-corrector (Welleck et al., 2022) | math word | finetune | corrector model | gen/corr feedback |
| Self-improvement (Huang et al., 2022a) | math word | finetune | self-assessment | CoT/consistency |
| Say-can (Ahn et al., 2022) | robot | model-generated | external model | greedy |
| Inner-monologue (Huang et al., 2022b) | robot | hand-written | various | greedy |
| Self-taught-reasoner (Zelikman et al., 2022) | math word | finetune | augmentation | greedy/feedback |
| Least-to-most (Zhou et al., 2022) | math word | hand-written | self-assessment | curriculum |
| Progressive-hint (Zheng et al., 2023) | math word | model-generated | self-assessment | stable prompt |
| Self-refine (Madaan et al., 2023) | math word | model-generated | self-assessment | greedy/feedback |
| Tree-of-thoughts (Yao et al., 2024) | puzzles | model-generated | self-assessment | BFS/DFS |
| Buffer-of-thoughts (Yang et al., 2024) | math word | thought template | self-assessment | buffer manager |
| Beam-search (Xie et al., 2024) | math word | model-generated | self-assessment | Beam Search |
| ReAct (Yao et al., 2022) | action | external knowledge | self-assessment | reinforcement learning |
| Reflexion (Shinn et al., 2024) | decision | model-generated | ext model | reinforcement learning |
| Voyager (Wang et al., 2023) | Minecraft | model-generated | Minecraft | reinforcement learning |
Domain
Many papers are applied to math word problems (natural language descriptions of math problems). Math problems were the original inspiration for the experiments with reasoning in LLMs. Other application domains include autonomous agents, robotic movement, generating computer programs, and playing computer games. We will discuss these in more detail with the individual approaches.
Taxonomy Table
Table 1 lists the papers of this survey. They are listed by the domain they work on, the type of prompt generation, the evaluation of the result, and the control method. The approaches in the table are grouped, divided by horizontal lines.
The first group, from Scratchpad to Self-ask, focuses on creating a prompt that generates the reasoning steps. The entries in the cells of this column are shown in bold, highlighting the focus of the approaches. The approaches in this group can be considered to be the start of the field of LLM-reasoning. The Chain-of-thought approach is especially an inspiration for many works. The prompts are often written “manually” by the researchers, the steps are encoded in one prompt, and step control is greedy. There is no specific evaluation of the steps, other than comparing results to the benchmark. The Scratchpad approach is special in that it uses supervised learning, not prompt-learning; the work showed that LLMs can be made to generate internal reasoning steps by supervised learning, paving the way for the later prompt-based papers.
The second group, from Self-verification to Self-taught-reasoner, focuses on evaluation of the reasoning steps in the prompt. This column is shown in bold in the table. The approaches in this group aim to improve the Chain-of-thought results by reducing the error accumulation that occurs when multiple steps are taken in a reasoning chain. A variety of step control methods is used by these approaches, which is discussed in more detail later. Note that not all approaches use natural language problems (often math word problems). For example, the subgroup of Codex to Program-aided-language focuses on formal languages. They generate code or math equations, typically in Python, to formalize the steps of the reasoning problem, or as result of the task. LLMs are quite good at code generation, and these approaches typically achieve good performance. The use of code also allows the approaches to call external programs such as interpreters and debuggers to evaluate the correctness of the reasoning steps that are generated.
There is also a special subgroup, Refiner to Self-improvement, that does not use prompt learning but finetuning. Here, new data is generated based on reasoning exemplars, which is then used to further train the model. The extra data is often generated as a separate dataset, sometimes called critic or corrector.
There are two approaches, Say-can and Inner-monologue, whose application domain is control of robot movement. Robotic movement is constrained by the laws of physics (both in the body of the robot as in aspects of its environment). The laws of physics are learned and used to ground the reasoning steps in reality (to reduce hallucination).
The third group, Least-to-most to Voyager, addresses step control (approaches shown in bold in this column). Whereas in the previous approaches the reasoning steps are written in a single, static, prompt, these approaches generate the steps in multiple, dynamic, prompts. This allows control of the space of reasoning steps. Various search control approaches are used, all in the form of an external algorithm that performs calls to the LLM with different prompts. The control methods range from simple greedy and depth-first search to elaborate beam search and reinforcement learning schemes.
In summary, we see a diverse array of methods that often achieve high performance in reasoning about their respective domains. To better understand the approaches, let us discuss the techniques in more detail, starting with the generation of steps.
5.1 Generation of Steps
Originally, LLMs performed poorly on math word problems (GSM8K (Cobbe et al., 2021)). Some different approaches were tried, for example scaling up the size of the LLM (Rae et al., 2021). The LLM architecture, based on transformers, is designed to produce a single token. When we prompt such an architecture to produce an answer, it does so. What we should do is prompt it to follow intermediate steps, answer those, and thus work towards the final answer, just as a student is taught to break down a complex problem into smaller steps. We should take the model by its hand and teach it to write down the intermediate steps, and combine the intermediate results (Nye et al., 2021).
This idea was used by Nye et al. (2021) in Scratchpads, a transformer model that performs multi-step computations by asking it to emit intermediate computation steps into a scratchpad. They train the model by supervised learning (not prompt-based in-context learning). Figure 2 shows an example. On experiments with addition, polynomial evaluation, and Python code execution, versions that produced the intermediate steps on a scratchpad performed considerably better than versions that did not.
If supervised learning can produce intermediate steps, would prompt-learning be able to do so too?
5.1.1 Hand-written Prompt
This question was studied by Wei et al. (2022b), amongst others. A basic way to instruct an LLM to generate steps by prompt-learning is to manually write a prompt for the large language model to follow the reasoning steps. They showed in their Chain-of-thought paper that with such a prompt the LLM follows such intermediate steps. When the LLM is prompted to rephrase information from the question as intermediate reasoning steps in its answer, the LLM performed much better than when it was prompted to answer a math problem directly, without reproducing the information from the question in its answer. The example from the Chain-of-thought paper is shown in Figure 3 Wei et al. (2022b). Performance figures were given in Section 3 on benchmarks.
The substantial performance improvement by Chain-of-thought has caused much excitement and has opened up further research on reasoning with LLMs. In the original Chain-of-thought paper the prompts were handwritten by the researchers for the individual types of problems, and evaluations are conducted with five different benchmarks (not by an LLM).333The Chain-of-thought idea is about prompt generation, not about the evaluation or the search control of the reasoning steps. Hence, in Table 1 Chain-of-thought is labeled as greedy without an evaluation. In a later work the prompts were generated automatically by the LLM (Zhang et al., 2022).
Kojima et al. (2022) go a step further. They show that the simple addition of a single text to the prompt (Let’s think step by step) significantly improves performance. Since this text does not contain problem-related elements, this can be considered as a form of zero-shot learning. Figure 4 compares the approaches. Experiments further show that with this addition to the prompt, significant performance gains are achieved on a diverse set of reasoning benchmarks, including arithmetic, symbolic, and logical reasoning.
5.1.2 Prompt using External Knowledge
Chain-of-thought shows that an LLM gives better answers to complex problems when it is guided to take individual steps. Prompts are written manually, from scratch, by the researchers.
We can use external information about the problem to improve the prompt. Press et al. (2022) study how subproblems are related to the main problem, which they call compositional reasoning. They study how often a model is able to answer the subproblems, but not the overall problem. This difference is called the compositionality gap. They find that in GPT-3, as model size increases, the compositionality gap does not decrease: the single-hop question-answering performance improves faster than the multi-hop performance. This shows that while more powerful models memorize and recall more factual knowledge, no improvement in their ability to perform compositional reasoning occurs. They find that the ability to reason does not depend on the size of the model.
Subsequently, a method called Self-ask is proposed, that asks elicitive follow-up questions (like Chain-of-thought, but with the follow up: prompt), see Figure 5. The model is then used to answer these follow-up questions. Self-ask can also use an external search engine to answer intermediate prompts, instead of the model. The model takes as input a compositional question which it decomposes. The initial subquestion is fed into the search engine, and the answer is processed by the model, which generates another subquestion, and so on, until it produces the final answer.
The approach performs a few percentage points better than vanilla Chain-of-thought on three benchmarks that were specifically designed for multi-hop questions.
5.1.3 Model-Generated Prompt
In addition to manually writing prompts or using external information, we can also try to let the LLM itself study the problem to write the best reasoning-prompt, a form of self-improvement. An example of this approach is Auto-chain-of-thought (Zhang et al., 2022). This approach builds on the observation by Kojima et al. (2022) that large language models are zero-shot reasoners. First, Auto-chain generates specific questions for a given dataset and partitions them into clusters. Then an external algorithm uses the model to generate examples that are sampled for diversity. The constructed demonstrations augment the in-context prompt. The automatically generated prompts are reported to perform as well or better than the hand-written Chain-of-thought prompts on ten benchmarks using GPT-3.
Fu et al. (2022) introduce Complexity-based prompting. Inspired by Chain-of-thought and Self-consistency, this work studies which prompts achieve the best results on math word and other reasoning problems. Their work specifically studies the impact of the complexity of the reasoning chain, and introduces a related reasoning approach (Complexity-based prompting). They find that prompts with the largest complexity (the most reasoning steps) perform best. Further, they find that outputs (answers) with the highest complexity are the best. Complexity-based prompting achieves high performance on three math reasoning benchmarks.
Another approach that uses model-generated prompts is Buffer-of-thoughts. We will discuss this approach in Section 5.3.3.
5.2 Evaluation of Steps
After discussing prompts for the generation of reasoning steps, the next stage in the reasoning pipeline (Section 2.3) is evaluation of the results of the steps, to reduce the error of multi-step reasoning chains.
We will start with approaches where the same model performs step-generation and step-evaluation.
5.2.1 Self-Assessment
When LLMs are prompted to perform reasoning steps, they perform a sequence of steps and predict multiple tokens. Performing a sequence of steps makes them sensitive to mistakes and vulnerable to error accumulation (Weng et al., 2022; Xiao et al., 2023a). Several methods have been developed to prevent error accumulation. One approach is to create a new model to separately evaluate the results. Shen et al. (2021) and Li et al. (2022b) train an external verifier to check results.
In contrast, Weng et al. (2022) propose an automated approach using evaluation by the same LLM, called Self-verification. They note that human reasoning also suffers from the problem of accumulating errors, and that in human reasoning we frequently revisit our thought process to verify the accuracy of our reasoning steps. Thus, they propose to apply such a backwards self-verification approach. The LLM is prompted to use the conclusion of the Chain-of-thought reasoning chain as a condition for solving the original problem and then compare the answer going back to the original question. The LLM is given variations of its own conclusion and is instructed to choose the one with the highest similarity to the original question. (Note that there can be feedback issues using an LLM to evaluate itself, for a discussion see Zheng et al. (2024).) Experiments are reported on GPT-3 (Chen et al., 2021) and on Instruct-GPT (Ouyang et al., 2022). The performance of Chain-of-thought was improved by a few percentage points on arithmetic and general reasoning tasks.
A popular related approach is called Self-consistency (Wang et al., 2022b). Self-consistency is a straightforward ensemble approach. Greedy single-path decoding is replaced by sampling diverse reasoning paths, evaluating them, and selecting the most consistent answer. Self-consistency asks the LLM to simply perform the same query multiple times, and takes the majority-vote of the answers. Self-consistency works since complex reasoning problems typically allow different reasoning paths that lead to the correct answer. Figure 6 summarizes the approach.
Self-consistency has been evaluated on arithmetic reasoning, commonsense reasoning and symbolic reasoning, on a variety of LLMs, including GPT-3 (Tay et al., 2022; Brown et al., 2020; Thoppilan et al., 2022; Chowdhery et al., 2023). Self-consistency improves the performance of Chain-of-thought typically by 10-20 percentage points, and has been used as a baseline in many of the other approaches in this survey. (Self-verification also reports that performance is improved when used in combination with Self-consistency (Wang et al., 2022b) and with Program-aided-language (Gao et al., 2023).)
5.2.2 Tool-based Validation
Another possibility to improve the accuracy of evaluating the reasoning steps is to switch from a natural to a formal language. The advantage of a formal language is that it is less ambiguous than a natural language. Examples are computer languages, such as Python, or mathematical equations. Using a formal language for reasoning is a popular approach, and we discuss seven papers. Many approaches generate the steps in Python, and the code can then be evaluated by a formal evaluator, such as a compiler, debugger, or interpreter.
LLMs have been quite successful in generating computer code from natural language prompts. Chen et al. (2021) introduced Codex, a GPT model that was trained on publicly available code in the repository GitHub. A production version of this work was introduced under the name GitHub Copilot. Codex is able to generate correct programs from descriptions in natural language, such as commentary strings. Figure 7 shows examples that are produced by Codex.
The work on Codex is used as a basis for further research on reasoning in LLMs.
Human programmers, when writing code, typically follow a cycle of writing some code, executing it to look for errors, and then using the feedback to improve the code. This same approach is followed in the Self-debugging work (Chen et al., 2023). Self-debugging teaches a large language model to debug its generated program code via few-shot demonstrations. It follows the same steps of (1) code generation, (2) code execution, and (3) code explanation (see Figure 8).
Self-debugging is able, without human feedback on the code’s correctness or error messages, to identify mistakes in the code that was generated by itself from investigating the execution results. Self-debugging can also provide an explanation of the generated code in natural language. It achieves strong performance on text-to-SQL generation, C++-to-Python transcoding, and text-to-Python generation.
Several works use self-debugging to generate working code tuned for solving specific problems automatically, without human feedback. Romera-Paredes et al. (2024) introduced FunSearch, an approach that integrates formal methods and LLMs to enhance mathematical reasoning and code generation. FunSearch is capable of producing functionally correct programs that adhere to specified requirements. It uses a genetic algorithm approach with multiple populations of candidate solutions (programs), which are automatically evaluated (using tools depending on the problem specification). In addition to the problem specification in the form of an evaluate function, also an initial program is given to the LLM in the first prompt. After evaluating a number of generated programs from the starting prompt, a new prompt using ‘best-shot prompting’ is created in an iterative fashion, combining a selection of sampled programs in a sorted list (ascending according to their evaluation score), and the LLM is requested to generate program . Another work leverages evolutionary computation methods to generate and optimize evolutionary algorithms (van Stein and Bäck, 2024). This approach, LLaMEA (Large Language Model Evolutionary Algorithm), utilizes LLMs to design and optimize evolutionary algorithms. The approach uses LLMs to generate initial algorithmic structures, which are then refined through mutation and selection. This enhances the efficiency of algorithm design, particularly in fields requiring innovative and adaptive solutions. A key difference between FunSearch and LLaMEA is that LLaMEA uses a sample-efficient elitism strategy by iteratively improving the best-so-far solution, requiring significantly fewer prompt evaluations than the large-population strategy proposed in FunSearch.
To improve prompt-based reasoning, Codex is used in an ensemble approach named MathPrompter (Imani et al., 2023). This approach generates multiple algebraic expressions or Python functions, which then solve the same math problem. The results are compared, just like in Self-consistency and Self-verification, raising the confidence level in the results. MathPrompter achieved state-of-the-art results on the MultiArith dataset (78.7% 92.5%), evaluated on GPT-3 175B.
Two other approaches that use a formal language are Program-of-thought (PoT) (Chen et al., 2022) and Program-aided-language (PAL) (Gao et al., 2023). Both approaches use the LLM to generate Python and then use a Python interpreter to evaluate the result. PoT and PAL are similar approaches. PoT uses benchmark-specific prompts; PAL uses generic prompts, and has been tested on more benchmarks and has been used in other approaches. Figure 9 illustrates the PAL approach.
When the evaluation of the reasoning steps is offloaded to the Python interpreter, decomposing the natural language problem into executable code-steps remains the only task for the LLM. (Earlier work in mathematical word problems, such as Ling et al. (2017), showed how to decompose a problem and reach an answer.) Gao et al. (2023) provide extensive experimental evidence about the synergy between the neural LLM and the symbolic interpreter. Experiments are performed over 13 mathematical, symbolic, and algorithmic reasoning tasks, achieving more accurate results than much larger models.
5.2.3 External Model Validation
We have seen many successful examples of prompt-based in-context reasoning and evaluation. We will now look at related reasoning approaches that follow a more traditional parameter learning approach. We describe three natural language approaches that follow this route. All approaches evaluate the output of the model and generate corrective data. That data is then added to the training pipeline, and the model is subsequently finetuned.
Finetuning
The Refiner approach (Paul et al., 2023) uses a generator model and a critic model to provide fine-grained feedback on reasoning errors. The generator generates multiple reasoning hypotheses, the critic evaluates results by randomly selecting a hypothesis for feedback. The generator model is finetuned based on its reasoning errors. A small supervised model is used to overcome the cold-start problem. Figure 10 shows an example of how the critic provides feedback to the generator.
The approach is reported to work well on math word problems and synthetic natural language reasoning.
Welleck et al. (2022) follow a similar approach in their Self-correction approach. The corrector is a separate model specialized in refining the outputs of the generator. Unlike Refiner, where the generator is finetuned based on the critic feedback, Self-correction finetunes the corrector to rectify errors in the hypotheses produced by the generator.
A third finetuning approach is Self-improvement, by Huang et al. (2022a). Here too the base model data is augmented by LLM-generated rationales, and then finetuned. Noteworthy in all three finetuning approaches is that LLMs are capable of improving themselves by training on their own generated output, and that stability problems inherent in feedback loops are overcome.
Dataset Augmentation
The final finetuning approach that we discuss uses dataset augmentation. An explicit intermediate reasoning is called a rationale. Rationale generation has been shown to be valuable for LLMs across diverse tasks such as mathematical and commonsense reasoning, code evaluation, social bias inference, and natural language inference (Zelikman et al., 2022). Zelikman et al. (2022) describe how reasoning steps are used to create rationales, that are then used to augment the dataset on which the model is finetuned. The approach is called Self-taught-reasoner. Figure 11 illustrates the approach.
In Self-taught-reasoner, an augmentation dataset is created by attempting to solve the original dataset using the current model’s rationale generation ability in each iteration. Next, the dataset is augmented using rationalizations, using ground-truth answers to problems the model failed to solve. Finally, the large language model is finetuned on the combined dataset.
Reasoning about Robot Behavior
In addition to math word problems, prompt-based reasoning has also been used to reason about robot behavior. Language models contain a large amount of information about the real world (Ahn et al., 2022). In theory, this should allow the model to exhibit realistic reasoning about robotic behavior. However, the models do not have knowledge about particular embodied aspects of a particular robot. If we could compare a Scratchpad-like list of intermediate reasoning steps with a list of possible movements of the robot in its environment, then we could prevent the model from suggesting impossible joint movements and actions, and prevent accidents.
Such an approach has been tried in the Say-can paper (Ahn et al., 2022). Say-can learns a value function (Kaelbling et al., 1996) of the behavior of a robot in an environment using temporal difference reinforcement learning Sutton (1988). This value function is then combined with prompt-based reasoning by the language model, to constrain it from suggesting impossible or harmful actions.
The goal of Say-can is to ground language in robotic affordances. In contrast to Scratchpad, which used supervised learning, the affordance model is learned interactively by reinforcement learning, and then applied using prompt-based learning by the LLM. The robot can act as the language model’s hands and eyes, while the language model has high-level semantic knowledge about the task. The LLM (Say) provides a task-grounding to find the actions to achieve the high-level goal. The learned affordance function (Can) provides a world-grounding to allow what is possible. Say-can is evaluated on 101 real-world robotic tasks, such as how to solve tasks in a kitchen environment (see Figure 12).
Where Say-can learns affordance as a separate function, another approach, Inner-monologue (Huang et al., 2022b) formulates robotic planning directly as part of the language prompt. This approach incorporates environmental information into the prompt, linguistically, as an inner monologue. As in Say-can, the information comes as feedback from different sources. Unlike Say-can, the information of physics and the world is inserted directly into the prompt.
Inner-monologue consists of many elements: it uses InstructGPT (Brown et al., 2020) for multi-step planning, scripted modules for object recognition, success detection, task-progress scene description, and language-conditioned pick-and-place primitives, similar to CLIPort (Shridhar et al., 2022). These elements generate textual descriptions that are used in prompt-based learning. Figure 13 gives an example of the working of Inner-monologue.
The language feedback that is thus generated significantly improves performance on three domains, such as simulated and real table top rearrangement tasks and manipulation tasks in a kitchen environment. There are many studies into robotic behavior. A recent approach related to Inner-monologue is Chain-of-tools, which proposes a plan-execute-observe pipeline to ground reasoning about tool behavior (Shi et al., 2024a, b).
This concludes our discussion of the second stage of the reasoning pipeline, evaluation of the reasoning steps.
5.3 Control of Steps
The third stage in the reasoning pipeline in Section 2.3 is reasoning control. This stage controls how many sub-steps are generated, and how deep into the future the reasoning chain is generated.
There are three main approaches: (1) greedy selection, which generates a step and then follows it, (2) ensemble strategy, which generates a set of possible next steps, and (3) a full tree-shaped search which generates multiple options for the step, and follows them multiple steps into the future, traversing a search tree with backtracking, controlling an exponential search space. We include reinforcement learning approaches, that interactively learn an optimal policy for such a reasoning space.
5.3.1 Greedy Selection
Most earlier works on prompt-based reasoning follow the greedy approach: generate a single prompt with a sequence of steps and follow them. Among the greedy reasoners are Chain-of-thought, Auto-CoT, and Zero-shot CoT. Inner Monologue and Say-Can also use greedy reasoning.
In Least-to-most prompting (Zhou et al., 2022), the key idea is to break down a complex problem into simpler subproblems and then solve these in sequence, explicitly encoding them in the prompt. It is related to Complexity-based prompting. In Least-to-most, finding the answer to each subproblem is facilitated by the answers to previously solved subproblems, as in a curriculum (Bengio et al., 2009). The authors find that on symbolic manipulation, compositional generalization, and math reasoning, the Least-to-most prompting is capable of generalizing to more difficult problems than those that are given in the prompts. Figure 14 illustrates the idea.
5.3.2 Ensemble Strategy
The second kind of reasoning control is based on an ensemble of (sequences of) reasoning steps. The ensemble approach is a well-known technique in machine learning to make a strong learner out of multiple weaker learners (Sagi and Rokach, 2018; Breiman, 2001). For most problems, multiple different options for the next step exist. When all or some of these are generated and evaluated, then the best result or the consensus result can be reported as the outcome of an ensemble of steps. Various approaches have been proposed.
We already mentioned Self-consistency (Wang et al., 2022b) and Self-verification (Weng et al., 2022) in Section 5.2.1. They are popular ensemble approaches to evaluate the results of reasoning steps in prompt learning. The greedy single-path decoding used in Chain-of-thought prompting is replaced by sampling a diverse set of reasoning paths, evaluating them, and selecting the most consistent answer.
In another domain Chain-of-experts builds on Chain-of-thought with a mixture of experts ensemble for complex combinatorial operations research problems (Xiao et al., 2023b). PAL and MathPrompter also use the ensemble approach. They generate multiple steps, which are evaluated and whose answer is combined, or the best step is chosen.
The ensemble approach is a popular approach in LLM-reasoning.
5.3.3 Reinforcement Learning
In the greedy approach, a single reasoning path is generated and traversed. In reasoning, often multiple valid reasoning steps are possible, but pursuing all possibilities over multiple reasoning steps may lead to an infeasible number of possibilities.
The third kind of reasoning control is to use a full-fledged controller that can traverse a tree, or even perform reinforcement learning to do so (Sutton and Barto, 2018; Kaelbling et al., 1996; Plaat, 2022). This group of control approaches enables the most elaborate control of the reasoning process, and is used by many works, as we will see. When decomposing the problem, multiple alternative steps are generated that can be searched multiple steps into the future. Then, backtracking can be performed, allowing alternative steps to be tried.
Where greedy and ensemble processes can be controlled with a prompt by the LLM, this third group is more complex, and an external algorithm is used to control the reasoning process. The external algorithms call the LLM as a subroutine prompting it to perform its tasks. This allows more complex reasoning control, but we are no longer performing prompt-based self-reasoning; control has been given to an algorithm that is external to the LLM and external to prompt-learning.
We start our discussion of control strategies with depth-first and breadth-first search, then go to beam search, and then to full reinforcement learning.
Breadth first search
A complex reasoning space can be traversed with a search algorithm. Tree-of-thoughts includes a search algorithm to dynamically follow different reasoning steps (Yao et al., 2024). When one reasoning path has been traversed, a search algorithm can backtrack, and try an alternative path. The paper describes both a breadth-first-search and a depth-first-search controller.
The evaluation part in Tree-of-thoughts is performed with a prompt by the LLM. Together, the trio of generation, evaluation, and control allow systematic exploration of the space of reasoning steps with look-ahead and backtracking. The authors compare their approach to Chain-of-thought and Self-consistency. Chain-of-thought builds a reasoning out of a path of thoughts, Self-consistency creates an ensemble of thoughts, and Tree-of-thoughts constructs a tree structure. Figure 15 illustrates the different reasoning structures.444A similarly named approach is Graph-of-thoughts (Besta et al., 2024). Graph-of-thoughts allows more general reasoning graphs, providing a formal framework, where the different elements can then be specified manually.
Another approach, Buffer-of-thoughts (Yang et al., 2024), goes a step further towards meta-reasoning. It introduces a meta-buffer that stores high-level thought-templates. These universal thought-templates are derived from a variety of tasks. Figure 16 compares the Buffer-of-thoughts approach to other approaches such as Chain-of-thought and Tree-of-thoughts. Buffer-of-thoughts outperforms other methods in puzzles such as Game of 24 and checkmating. Thought templates are related to metacognition (thinking about thinking), which is further discussed in Section 6.2.3.
Beam search
A related search method is Beam-search. Beam-search-for-reasoning (Xie et al., 2024) focuses on control of the space of possible reasoning paths. In some reasoning problems, this space can be very large. Beam-search solves this challenge by searching only a promising part of this space. It uses self-evaluation to control exploration and to evaluate (decode) reasoning steps. Figure 17 shows how Beam-search self-evaluation is used in multi-step reasoning.
Beam search uses Program-aided-language models for math word problems (Gao et al., 2023). Using a Codex backbone (Chen et al., 2021), it surpasses the few-shot baselines by 6.34%, 9.56%, and 5.46% on the GSM8K, AQuA, and StrategyQA benchmarks, respectively.
Reinforcement learning
Reinforcement learning (RL) methods are another step in the sophistication of optimization algorithms. RL learns by interactive sampling, improving its policy based on rewards from the environment (Sutton and Barto, 2018). To use reinforcement learning, the reasoning problem must be formulated as a Markov Decision Process: the agent-algorithm creates a prompt (an action), to sample a step () and get an answer (state, reward) from the environment-model (see Figure 18). The answer can then be used to improve the prompt (next action), just like reinforcement learning uses rewards to improve its policy of best actions for each state. The approaches that use reinforcement learning do so in the form of an external algorithm. No prompt has been created that performs RL by itself.
Progressive-hint-prompting (PHP) uses reinforcement learning to interactively improve prompts (Zheng et al., 2023). Figure 19 illustrates the approach.
PHP is an external algorithm that calls the LLM with dynamic prompts, using previously generated answers as hints to progressively prompt the LLM toward the correct answers. It works as follows: (1) given a question (prompt), the LLM provides a base answer, and (2) by combining the question and answer, the LLM is queried and we obtain a subsequent answer. We (3) repeat operation (2) until the answer becomes stable, like a regular policy-optimizing reinforcement learning algorithm. The authors have combined PHP with Chain-of-thought and with Self-consistency. Using GPT-4, state-of-the-art performance was achieved in grade school math questions (95%), simple math word problems (91%) and algebraic question answering (79%).
Another approach that is motivated by improving answers from feedback, is Self-refine (Madaan et al., 2023). In this method, initial outputs from LLMs are improved through iterative feedback and refinement. Like PHP, the LLM generates an initial output and provides feedback for its answer, using it to refine itself, iteratively. Figures 20 and 21 illustrate the approach.
Self-refine prompts the LLM in three ways: (1) for initial generation, (2) for feedback, and (3) for refinement. Note that Self-refine follows a greedy reasoning chain, learning from feedback. Self-refine has been used with GPT-3.5 and GPT-4 as base LLMs, and has been benchmarked on dialogue response generation (Askari et al., 2024), code optimization, code readability improvement, math reasoning, sentiment reversal, acronym generation, and constrained generation, showing substantial improvements over the base models.
Another approach that combines reinforcement learning and LLMs is ReAct (Yao et al., 2022). Most works so far have focused on reasoning by the LLM, not on actions by an agent. A key element of reinforcement learning is that it learns a policy for an environment. The goal of ReAct is to combine progress in reasoning with action plan generation. (Or, to put it differently, most approaches use RL to improve LLM-reasoning, ReAct uses LLMs to improve RL agent policies.) ReAct uses Chain-of-thought prompt-learning as part of an RL framework that also uses external knowledge sources (Wikipedia) and finetuning, for error reduction, grounding, and for reducing hallucination. The framework allows hand-written prompts. Figure 22 shows four different prompting strategies.
On two interactive decision making benchmarks (ALFWorld and WebShop), ReAct outperforms imitation and reinforcement learning methods by an absolute success rate of 34% and 10% respectively, while being prompted with one or two in-context examples.
The ReAct work has been developed further. Reflexion (Shinn et al., 2024) is built on top of ReAct. The goal is to create AI agents that learn by reflecting on failures and enhancing their results, much like humans do. Reflexion uses three language models: actor, evaluator, and reflector. It works as follows: (1) an actor generates text and actions, (2) an evaluator model scores the outputs produced by the actor, and (3) a self-reflection model generates verbal reinforcement cues to assist the actor to self-improve (see Figure 23). For the actor, Chain-of-thought (Wei et al., 2022b) and ReAct (Yao et al., 2022) can be used. Reflexion is evaluated on decision-making, reasoning, and coding tasks. Improvements of 10-20 percentage points are reported. Figure 24 shows three different prompting applications.
To conclude this overview of reinforcement learning approaches, we discuss an application in the games domain. Voyager (Wang et al., 2023) is an agent for the game of Minecraft that uses an iterative prompting mechanism that generates code for embodied control. The mechanism includes Self-verification (Shinn et al., 2024). The agent has a skill library and an automatic curriculum to maximize exploration. Voyager interacts with GPT-4 through prompts. The goal of Voyager’s prompts is to discover as many diverse items in Minecraft as possible, a form of novelty search (Eysenbach et al., 2018). Voyager performs well, it shows in-context lifelong learning capability and reaches high scores by acquiring many tools (see Figure 25).
6 Discussion
We have reviewed approaches for prompt-based reasoning by LLMs, highlighting techniques that have achieved a breakthrough in reasoning performance. It is time for reflection on limitations in the approaches, suggesting promising areas of future work. First we discuss issues concerning hallucination, faithful reasoning, and scaling. Then we discuss what LLMs can and cannot do. Then, we highlight connections with sequential decision processes and metacognition, and end with a research agenda.
6.1 Hallucination, Faithfulness and Scaling
Most works on reasoning in LLMs are experimental in nature. The success of in-context learning and Chain-of-thought reasoning is attracting the attention of work providing deeper insight into the reasoning processes in language models.
Saparov and He (2022) introduce a synthetic question/answer dataset designed to evaluate the reasoning abilities of LLMs. The work showed that LLMs are capable of reasoning to a certain degree, but that Chain-of-thought struggles with proof trees with a wide branching factor. In another study, Wang et al. (2022a) also aim to increase our understanding of how Chain-of-thought works. The authors find that it continues to work even with invalid steps in the reasoning chain. They also find that the order of the reasoning steps is important for good results. Prompts should be relevant to the question, and coherent (steps should be in the correct order). Jin et al. (2024) study the impact of reasoning step length on LLMs, finding a strong positive correlation between the length of the prompt and reasoning abilities.
These works highlight ways in which LLM-reasoning can see things that are not there. Next, we discuss works on failure modes of the Chain-of-thought approach, studying whether the reasoning of the LLM is faithful, or that it gives the right answer for the wrong reason.
6.1.1 Faithfulness
Chain-of-thought and other approaches prompt a language model to take certain steps to solve the problem that the prompt specifies. One can ask the question, whether those steps are indeed the steps that the model has followed (faithful reasoning) or whether it took another road to arrive at the correct answer (unfaithful reasoning). A few studies measure the faithfulness of reasoning by LLMs. Lanham et al. (2023) notes that just like organic reasoners, a model’s reasoning may be post-hoc, it may be constructed after a certain conclusion has been found. By deliberately adding mistakes to the chain of thought, the authors measure the faithfulness of the model. They find a wide variation of post-hoc reasoning, with a tendency of larger models to be less faithful. Like regular LLMs, when not properly grounded, (Chain-of-thought) reasoning suffers from hallucination.
Another study adds deliberate bias to the prompt. For example, in a multiple-choice setting, they always make answer (a) the correct answer (Turpin et al., 2024). They find that a bias towards wrong answers can cause significant drops in accuracy, and that models frequently generate Chain-of-though explanations rationalizing wrong answers. The authors further note that, insofar as language models are trained on human-written explanations, that explanations may be incomplete or wrong. Human explanations may omit crucial steps of the causal chain, may provide an unfaithful account of the human reasoning process, or may be aimed at convincing others, instead of providing the true causes of a decision.
To address issues of faithfulness, Lyu et al. (2023) propose Faithful-chain-of-thought. This approach involves two stages. First, the natural language query is translated into a formal symbolic language. Second, the problem-solving stage processes the formal language, and can explain the reasoning steps it has thus taken. For the symbolic language, Python, Datalog, or PDDL is suggested. Faithfulness studies tell us more about how models reason. Further surveys on this topic are Mondorf and Plank (2024); Chuang et al. (2024); Luo et al. (2023); Paul et al. (2024),
6.1.2 Scaling
The emergent abilities of LLMs have prompted research into the nature of scaling and reasoning with LLMs, and, specifically, how reasoning capabilities can be transferred to smaller language models. Scaling laws of LLMs are an active area of study, see for example Kaplan et al. (2020); Henighan et al. (2020); Hoffmann et al. (2022). Given the computational cost of LLMs, there is much interest in transferring knowledge to small language models. Comprehensive surveys on knowledge distillation are Xu et al. (2024); Gu et al. (2023). For reasoning specifically, Magister et al. (2022) have studied reasoning in small language models, using a student model that learns from a teacher model, by finetuning. Another study related to Self-taught-reasoner (Li et al., 2022a) focuses on explanation in small language models, achieving similar results.
Other works focus on prompt distillation for retrieval Dai et al. (2022), recommendation (Li et al., 2023), distillation to embodied agents of Chain-of-thought reasoning (Choi et al., ), and distillation of LLM graph reasoning (Zhang et al., 2024). Distillation of reasoning to smaller models can work surprisingly well in situations with more explicit instructions. Distillation is also proposed for bringing results of System 2 reasoning to System 1 Yu et al. (2024), which brings us to the topic of metacognition (see Section 6.2.3).
6.2 Limitations: What LLMs Can and Cannot do
The capabilities of LLMs are impressive. LLMs can be seen as large text-based surrogate models of the world (or the world how we describe it on the internet), and thus allow us to reason in a way that we can understand about a large variety of contexts and problems. Reasoning tasks, such as math word problems, were one of the capabilities that LLMs could not achieve, until recently. Let us look more closely at what language models can and cannot do.
6.2.1 What Can LLMs Do?
With the right prompt LLMs are able to solve many of the problems in reasoning grade school math word benchmarks. Prompt-based learning is able to perform reasoning tasks such as math word problems, robotic movement, and Python code generation, at inference time, without expensive parameter training.
We note that a simple taxonomy of generate-evaluate-control is able to describe the structure of the current LLM reasoning literature well. Furthermore, the accuracy of the reasoning chains can be improved with ensemble methods, or self-verification. Hallucination can be reduced by grounding the model with external models, such as for robotic affordances, and information retrieval from search engines and Wikipedia. Going a step further, using external control algorithms (such as search or RL) as scaffolding, dynamic prompts can use the LLMs to perform complex and interactive reasoning patterns.
Note that the reasoning control is now two layers away from the core LLM: an external control algorithm, on top of in-context-learning, dynamically generating prompts for the LLM. This is reasoning with prompts with LLMs, not by.
At this point, it is interesting to note the confluence of the two schools of classical artificial intelligence (AI), symbolic and connectionist.555Reasoning and planning have been studied since the start of artificial intelligence, starting with logic and reasoning (Newell and Simon, 1961), search algorithms in puzzles and board games (Korf, 1999; Plaat, 2020), robot planning (Fikes and Nilsson, 1971), classical machine learning such as decision trees and support vector machines (Flach, 2012; Breiman, 2001; Cortes and Vapnik, 1995), through knowledge representation and the semantic web (Van Harmelen et al., 2008). Ever since the success of the connectionist approach LeCun et al. (2015); Goodfellow et al. (2016) (deep learning, including LLMs) researchers have tried to join the two approaches. Search and reinforcement learning are rooted in the symbolic AI tradition, while LLMs are rooted in the connectionist tradition. The literature in this survey combines the two traditions. High performance reasoning is created with a (symbolic) searcher/learner on top of a (connectionist) LLM. In other fields similar combinations can be seen (for example, AlphaFold Bryant et al. (2022); Jumper et al. (2021) and retrosynthesis of molecules Segler et al. (2018)). The LLM helps ground symbolic reasoning methods in language; symbolic methods help create prompts that let the LLM perform reasoning. How the two traditions will continue to improve eachother, we will see in further research.
We note that benchmarks such as GSM8K have been central for the progress in the field, and that while reasoning started with math word problems, the field has extended to robotics, autonomous agents, games, and most emphatically computer code. Formal languages play an important role in the intermediate multi-step reasoning chains.
A side effect from the work on reasoning is the emergence of a new few-shot learning approach for sequential decision-making processes (SDP)(Littman, 1996). Traditionally these processes are solved with reinforcement learning (such as DQN Mnih et al. (2015), PPO (Schulman et al., 2017) and SAC Haarnoja et al. (2018)), achieving good results, but suffering from high sample complexity for larger problems Plaat et al. (2023). The emergence of few-shot in-context learning for solving SDPs opens a research avenue to find out what SDPs few-shot prompt-learning will be able to solve.
6.2.2 What Can LLMs Not Do?
Now that grade school math word problems are largely solvable, harder reasoning benchmarks in other domains are appearing (Ahn et al., 2024). Another line of research argues that LLMs cannot reason, providing examples where LLMs fail, and discussing potential reasons. Berglund et al. (2023) show that LLMs can fail to generalize in surprising ways. They provide the example that if a model is trained to report that ”Valentina Tereshkova was the first woman to travel to space”, it will not automatically be able to answer the question, ”Who was the first woman to travel to space?” pointing to a lack in semantic understanding of LLMs. Other work suggests that results are less generalizable and transferable than often assumed, showing how base-10 arithmetic skills do not transfer to base-9 arithmetic problems Wu et al. (2024). The question which problems LLMs can and cannot solve will continue to motivate researchers.
Other works study the dangers of the size of LLMs. Bender et al. (2021) mention the environmental risks associated with the large computational training demands, as well as the difficulty of understanding the training data, for example in the context of bias. Furthermore, there are ethical, legal, and copyright concerns regarding the data that LLMs are trained on. Finally, to prevent putting too much trust in the outcome of LLMs, we should understand their failure modes better, such as the well-publicized problems of hallucination (inventing facts that look right but are not).
Most of the reasoning capabilities exhibited by LLMs are due to the great representational powers of the transformer architecture, and how in-context learning is able to harness them. Prompt engineering and prompt control play a crucial role in the kind of reasoning that we have seen in the papers. Models can be instructed to write their own reasoning prompts; however, such Auto-GPT or Auto-CoT prompts need evaluation, verification, and grounding in the real world, to prevent degeneration into a hallucinatory world of their own. Models can also be instructed to interact with the world, and become the tool of external scaffolding that evaluates, controls and improves the prompts. Some of what we experience as reasoning by the LLM, is controlled by the prompt or the scaffolding algorithm. It is an open question if prompt learning is able get the LLM to create a prompt to exhibit non-trivial reasoning by itself.
From the symbolic planning field there is also a critical view on the reasoning and planning abilities of LLMs (Valmeekam et al., 2023) giving examples of planning failures. They argue that LLMs can be used instead to improve heuristic elements of traditional planners, such as PDDL (Kambhampati et al., 2024), to strengthen traditional symbolic planning approaches.
Some of the names of the approaches surveyed in this paper are suggestive of self-awareness and self-reflective capabilities. True self-reflection, or metacognition, is still largely outside the capabilities of current LLMs. LLMs can be prompted to reason, to take small steps, to self-evaluate, and their search process can be controlled by an external algorithm. The self-reflective type of “intelligence” is written into the prompt by the prompt engineer or the interactive algorithm. We are unaware of any LLM that has been made to reflect on, or even control, its reasoning processes, controlling how many reasoning steps it should take, or limiting its reasoning once the answer had become good enough. True self-reflection remains future work, although some steps have been taken, as we will discuss next.
6.2.3 Reasoning towards Metacognition
Human thought exhibits the ability to reason about self, we are able to think about our own thinking processes. Metacognition studies these topics (Veenman et al., 2006). Prompted by the success of Chain-of-thought and the works that we have surveyed, metacognition has also been studied in the context of LLMs (Toy et al., 2024).
Many reasoning approaches highlight self-reflective aspects in their names and in how they work. The prompts that prompt the models to reason are being improved with the outcome of the reasoning process, and in Buffer-of-thoughts thought-templates are used that are derived from other reasoning processes. Wang and Zhao (2023) study Metacognitive-prompting. Inspired by Chain-of-thought and Self-consistency, they create manually designed prompts to increase the understanding of language models. Figure 26 illustrates the relation between metacognitive human thought processes and metacognitive LLM prompting.
Another work, again inspired by Chain-of-thought and Self-consistency, connects psychology and LLMs. Didolkar et al. (2024) study metacognitive capabilities of LLMs in mathematical problem solving, both on GSM8K and on the harder MATH problems (Hendrycks et al., 2021). First, the model is prompted to find a skill name for each problem instance in the dataset. For 7000 instances of GSM8K, 500 skill names were found by the model. Next, these 500 names are clustered down to 22 skills. They find that by using the names of these 22 skills in Chain-of-thought-like prompts, more problems are solved than with standard Chain-of-Thought/Self-consistency/PAL prompts. Examples of the 22 skill names are multiplication-and-addition, basic-arithmetic, subtraction, and algebra. Interestingly, the authors find that the skill exemplar repository that is trained on a strong model (GPT-4), also down-translates to a weak model (GPT-3). The performance of the weak model benefits from the skill-name-enhanced prompts.
The connection between reasoning in LLMs and full-blown metacognitive reasoning is in its early stages. Exciting future research may appear.
6.3 Research Agenda
At the end of this discussion, we present promising topics for future work. Reasoning with LLMs is an active field of research. It brings together elements of symbolic reasoning, connectionism, natural language, autonomous agents, and affective reasoning (Broekens et al., 2023) with the promise of artificial general intelligence.
For the future, the surveyed works point in the following directions. First we discuss topics for the field of LLM-reasoning itself, then we discuss more general machine learning topics that are important for progress in LLM-reasoning, and finally we discuss more longer term, fundamental topics.
Specific research topics for reasoning with LLMs are:
Control and prompt-learning—Search control beyond greedy search is implemented as an external algorithm. Is it possible to incorporate all stages of the reasoning pipeline into an interactive prompt? Can we make a prompt that performs dynamic search-like step control without external scaffolding?
Code—Progress in reasoning using formal languages and computer code has been quite promising. GitHub Copilot is a success. Further integration of LLM-reasoning with software engineering tools is a promising area of research that can have a large practical impact on how software is written.
Grounding—Reasoning in LLMs has been successfully applied in autonomous agents, robotics, and games. A challenge is the grounding of the reasoning process in the environment. How can we help LLMs to actively find new information when the reasoning outcome is uncertain? Is retrieval augmented generation the future? Is the future of the reasoning-LLM a search engine (Verberne, 2024)?
Generic topics in machine learning that also influence prompt-based reasoning research are:
Benchmarks—Progress in LLMs is governed by the availability of the right benchmarks. The current favorite is GSM8K, for grade school math. As the field progresses, other benchmarks will become prevalent: benchmarks with more difficult tasks, and benchmarks for other applications in autonomous agents and robotics.
Faithfulness—Our theoretical understanding of prompt-based reasoning with LLMs is incomplete. The research on faithfulness highlights one example of our lack of understanding. In general, more insight into the working of multi-step in-context learning in LLMs is dearly needed.
Small language models—Efficiency is an important element for wide adoption of language models. Important topics are distillation of reasoning to small language models and an understanding of scaling laws.
Few-shot Reinforcement Learning—Small reasoning problems can be solved with few-shot in-context learning. Can we solve larger sequential decision processes, reducing the sample complexity in reinforcement learning?
For longer term future work, the following more fundamental questions are important:
Symbolic and Connectionist Computation—How can we further improve LLM-reasoning: how can LLMs benefit from symbolic reasoning prompts and how can LLMs help ground symbolic reasoning in language?
Metacognition—Much of the research into reasoning guides the model how it should solve a problem. Is it helpful to introduce named concepts for different kinds of reasoning? Can the model find these concepts by itself? Making the LLM “think” step by step is a first step towards influencing the model’s own “thought” processes. The first works on LLM metacognition have appeared, and artificial general intelligence will pursue this further.
7 Conclusion
Prompt-based in-context learning is an efficient machine learning method, requiring no parameter updates to the LLM. While achieving good performance on language tasks (System 1), performance on reasoning tasks (System 2) was lacking. Reasoning tasks, such as math word problems, are typically solved in step-by-step fashion. Recently prompts have been developed that guide an LLM to “think step by step” (Chain-of-thought), and to evaluate and verify the step results. The performance of reasoning with LLMs has improved greatly. Together, the surveyed methods allow the LLM to follow high-quality multi-step reasoning chains. Python code or other formal languages have been used successfully to reduce the error in reasoning steps. Also, in the field of autonomous agents and robotic action, good performance has been achieved by grounding reasoning answers in the environment and the physical constraints of robotic movement.
For complex reasoning tasks a large number of reasoning steps may be generated. To control the size of the reasoning space interactively, external scaffolding algorithms can be used. Often, variations on search algorithms or reinforcement learning are used. The symbolic and connectionist AI traditions come together in reasoning prompts and search algorithms that help LLM neural networks solve natural language math word and related problems.
Among the most popular reasoning benchmarks in this survey is GSM8K, which contains 8500 grade school math word problems. With LLMs such as GPT-3, reasoning approaches show an improvement of 20-50% points over standard prompting methods. For further progress in the field, the development of other challenging benchmarks is important.
The field of reasoning with LLMs is quite new, and theoretical understanding is lacking in important areas, such as faithful reasoning (models may sometimes find the right answer for the wrong reason). Although prompt-based learning allows few-shot learning, the computational needs of LLMs pretraining and finetuning are still high, hence the interest in small language models. Reasoning skills that work in large models can often be transferred to small models.
Human thought is capable of metacognition, we can think about our thinking process. Many of the names of the approaches in this survey suggest a link to metacognition (Reflexion, Self-refine, Self-improvement, Inner-monologue). The first preliminary experiments of language models that reason about their reasoning skills have appeared.
LLM-reasoning is an active field of research, with connections to artificial general intelligence. The field has shown great progress. Based on current limitations and open questions we provide a research agenda highlighting opportunities for further progress in harder reasoning problems, metacognition, and small language models, amongst others.

### v2

@@
-
-Abstract
-Scaling up language models to billions of parameters has opened up possibilities for in-context learning, allowing instruction tuning and few-shot learning on tasks that the model was not specifically trained for.
-This has achieved breakthrough performance on language tasks such as translation, summarization, and question-answering.
-Furthermore, in addition to these associative “System 1” tasks, recent advances in Chain-of-thought prompt learning have demonstrated strong “System 2” reasoning abilities, answering a question in the field of artificial general intelligence whether LLMs can reason.
-The field started with the question whether LLMs can solve grade school math word problems.
-This paper reviews the rapidly expanding field of prompt-based reasoning with LLMs.
-Our taxonomy identifies different ways to generate, evaluate, and control multi-step reasoning.
+Multi-Step Reasoning with Large Language Models, a Survey
+Abstract.
+Language models with billions of parameters exhibit in-context learning abilities, enabling few-shot learning on tasks that the model was not specifically trained for.
+Traditional models achieve breakthrough performance on language tasks, but do not perform well on basic reasoning benchmarks.
+However, a new in-context learning approach, Chain-of-thought, has demonstrated strong multi-step reasoning abilities on these benchmarks.
+The research on LLM reasoning abilities started with the question whether LLMs can solve grade school math word problems, and has expanded to other tasks in the past few years.
+This paper reviews the field of multi-step reasoning with LLMs.
+We propose a taxonomy that identifies different ways to generate, evaluate, and control multi-step reasoning.
 We provide an in-depth coverage of core approaches and open problems, and we propose a research agenda for the near future.
-Finally, we highlight the relation between reasoning and prompt-based learning, and we discuss the relation between reasoning, sequential decision processes, and reinforcement learning.
-We find that self-improvement, self-reflection, and some metacognitive abilities of the reasoning processes are possible through the judicious use of prompts.
-True self-improvement and self-reasoning, to go from reasoning with LLMs to reasoning by LLMs, remains future work.
-1 Introduction
-Transformer-based Large Language Models (LLMs) that are trained on large datasets have achieved breakthrough performance at next token prediction (Vaswani et al., 2017; Radford et al., 2019; Wei et al., 2022a); they are very good at natural language understanding (GLUE, SQUAD, Xsum) (Wang et al., 2018, 2019; Rajpurkar et al., 2016; Narayan et al., 2018), translation (Kocmi et al., 2022; Papineni et al., 2002; Sennrich et al., 2015), question answering (Tan et al., 2023), and other System 1 tasks (Kahneman, 2011).111In his book Thinking, fast and slow, a bestseller on human psychology, Daniel Kahneman described System 1 thinking as a near-instantaneous process; it happens automatically, intuitively, and with little effort.
-It is driven by instinct and experiences.
-System 2 thinking is slower and requires more effort.
-It is conscious and logical.
-The automatic operations of System 1 generate surprisingly complex patterns of ideas, but only the slower System 2 can construct thoughts in an orderly series of steps.
-In the LLM literature the terms are often used as shorthand to distinguish single-step associative tasks, from multi-step reasoning tasks, despite the fact that language tasks such as question answering and translation may require some “slow” thinking.
-The success of ChatGPT (Ouyang et al., 2022) has taken the world by storm.
-Transformer-based generative language models whose size is beyond hundreds of billions parameters are not only very good at language generation, they also enable new type of machine learning, called in-context learning (Brown et al., 2020).
-In-context learning, also known as prompt-based learning, occurs only in LLMs beyond a certain size (hundreds of billions of parameters) that are sufficiently rich (Wei et al., 2022a).
-In-context learning is inference time, prompt-based, few-shot learning, where model parameters are not trained or fine-tuned.
-System 1 tasks, such as associative language tasks, are easily solved by LLMs with prompt-based learning, as the many school children around the world that use ChatGPT daily can attest.
-(Although the problems are too often not solved correctly, just fluently, when the model’s association powers lead to hallucination (Huang et al., 2023).) On the other hand, System 2 tasks, such as grade school math word problems, are more difficult for LLMs(Cobbe et al., 2021).
-To solve math word problems we need to break down the problem in multiple reasoning steps.
-Spurred-on by the impressive performance on System 1 tasks, much research has focused on understanding the reason for the poor performance of LLMs on System 2 tasks, and how it can be improved.
-Among this research, the Chain-of-thought experiment (Wei et al., 2022b) stands out.
-This work, and subsequently Kojima et al.
-(2022), showed that adding a simple instruction to the prompts, Let’s think step by step, can provoke an LLM to perform the required intermediate reasoning steps, achieving a surprising jump in performance.
-The Chain-of-thought paper is a breakthrough in the field of reasoning with LLMs.
-Much exciting work has been published that builds on this work.
-Grade school math word problems started the research into LLM-reasoning, with the GSM8K benchmark (Cobbe et al., 2021).
-In our survey we discuss papers based on this benchmark, and directly-related follow up work on reasoning.
-We focus on prompt-based approaches.
-We survey the recent literature using a straightforward taxonomy.
-Although the field has only recently started, the jump in performance on reasoning has excited artificial intelligence and society alike.
-We provide a research agenda with opportunities for future research.
-At the end of this survey, we also discuss connections to other fields, such as self-reflection, metacognition (or thinking about thinking, see for example Dunlosky and Metcalfe (2008)), and the motivation towards artificial general intelligence.
-Our contributions are:
-A survey of relevant approaches in prompt-based reasoning (grade school math word problems and closely related domains) in large language models, including a research agenda.
-A taxonomy based on regular reasoning literature (step generation, step evaluation, and control of reasoning steps).
+We find that multi-step reasoning approaches have progressed beyond math word problems, and can now successfully solve challenges in logic, combinatorial games, and robotics, sometimes by first generating code that is then executed by external tools.
+Many studies in multi-step methods are using reinforcement learning for finetuning, external optimization loops, in context reinforcement learning, and self-reflection.
+1. Introduction
+Transformer-based Large Language Models (LLMs) that are trained on large datasets have achieved breakthrough performance at text generation tasks that directly build on next token prediction (Vaswani et al., 2017; Radford et al., 2019; Wei et al., 2022a); they are good at natural language understanding (GLUE, SQUAD, Xsum) (Wang et al., 2018, 2019; Rajpurkar et al., 2016; Narayan et al., 2018), translation (Kocmi et al., 2022; Papineni et al., 2002; Sennrich et al., 2016), question answering (Tan et al., 2023), and other language generation tasks.
+The success of models such as ChatGPT (Ouyang et al., 2022) is impressive.
+Transformer-based generative language models whose size is beyond hundreds of billions parameters are not only good at language generation, they also enable a new type of machine learning, called in-context learning (Brown et al., 2020).
+In-context learning, also known as prompt-based learning, is an emergent ability that occurs in LLMs beyond a certain size (hundreds of billions of parameters—less, with judicious prompting) that have been finetuned for conversational responses (Wei et al., 2022a).
+In-context learning is inference-time, prompt-based, few-shot learning with instructions.
+As opposed to finetuning, model parameters are not adapted by in-context learning.
+Language generation tasks are solved well by LLMs with prompt-based learning.
+On the other hand tasks that require reasoning, such as grade school math word problems, are more difficult for LLMs (Cobbe et al., 2021).
+Spurred-on by the impressive performance on language tasks, much research has focused on understanding the reason for the poor performance of LLMs on reasoning tasks, and how it can be improved.
+Among this research, the Chain-of-thought paper by Wei et al.
+(2022b) stands out.
+This work, and later work by Kojima et al.
+(2022), showed that adding a simple instruction to the prompt, Let’s think step by step, can provoke an LLM to perform the required intermediate reasoning steps to answer difficult problems in a multi-step approach.
+Subsequently, performance on math word benchmarks has increased markedly.
+Much of the performance increase of models such as OpenAI o1, o3 (Huang et al., 2024) and DeepSeek R1 (Shao et al., 2024; Guo et al., 2025) is attributed to multi-step reasoning methods as reviewed here, with reinforcement learning playing an increasingly important role, both for in-context learning and finetuning (Xu et al., 2025; Wu et al., 2024a; Du et al., 2025; Chu et al., 2025).
+The line of research into multi-step LLM-reasoning was started with grade school math word problems, with the GSM8K benchmark (Cobbe et al., 2021).
+Soon, other reasoning domains were included, such as reasoning about advanced math problems, computer code, robotic movement, and games.
+Many works have been published that build on Chain-of-thought (Chu et al., 2024b).
+In this paper, we survey the literature using a straightforward taxonomy.
+We discuss papers based on several reasoning benchmarks, as well as directly-related follow up work.
+Having started with basic math word problems, multi-step LLM reasoning approaches now perform logic reasoning, planning, combinatorial games, and robotic actions, amongst others.
+Some approaches do so by prompting the LLM to generate code that is then interpreted by external tools.
+Many studies in multi-step methods are using reinforcement learning, for finetuning, in context reinforcement learning, external optimization loops, and self-reflection.
+The main contributions of this paper are: (1) we provide a survey of relevant approaches in multi-step reasoning with LLMs, (2) we propose a taxonomy based on the reasoning literature (step generation, step evaluation, and control of reasoning steps), and (3) we formulate a research agenda for reasoning with LLMs.
 This survey is organized as follows.
-Section 2 summarizes the most relevant developments in LLMs, including in-context learning.
-Of great importance are the benchmarks that are used in this field.
-We discuss these in Section 3, followed by our method for scoping and selecting of papers in Section 4.
-Next, in Section 5 we provide a taxonomy of the field, where we discuss the approaches in detail.
+Section 2.1 provides background information on the most relevant developments in LLMs, including in-context learning and finetuning.
+Of great importance are the benchmarks that are used in this field (Section 2.4).
+Next, in Section 3 we provide a taxonomy of the field, where we discuss the main approaches in detail.
 Then, in Section 4 we discuss our findings in a broader perspective.
-We also discuss the relation between reasoning and work on self-reflection and metacognition.
+We also discuss the relation between multi-step reasoning and work on self-reflection and metacognition.
 This section concludes with an agenda for future research.
 Finally, Section 5 concludes the survey.
-2 Background: Reasoning with LLMs
-Before we dive into the works on reasoning, we review some background terminology on LLMs.
+2. Background, Scope and Selection of Papers
+Reasoning has a long and active history in AI, in logical inference, and in other fields, such as commonsense and analogical reasoning (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990).
+Indeed, some of the early criticism on LLMs was that they could not reason, and that they made obvious and basic reasoning errors, showing a lack of understanding.
+Berglund et al.
+(2024) provide the example of a model that is trained to report that “Valentina Tereshkova was the first woman to travel to space,” but is not able to answer the question, “Who was the first woman to travel to space?” pointing indeed to a lack of semantic understanding.
+Other work suggests that results are less generalizable and transferable than often assumed, showing how base-10 arithmetic skills do not transfer to base-9 arithmetic problems (Wu et al., 2024b).
+Different kinds of reasoning have played an important role in LLM research.
+2.1. Scope
+The interest in Chain-of-though has stimulated more research into language and logic.
+Logic-LM combines LLMs with symbolic logic solvers in an agentic way, where the symbolic solver is called by the LLM as a tool (Pan et al., 2023).
+Xu et al.
+(2024a) use a similar approach.
+Relevant logic-reasoning benchmarks are PrOntoQA (Saparov and He, 2023), ProofWriter (Tafjord et al., 2020), FOLIO (Han et al., 2022), LogicalDeduction (Srivastava et al., 2022), and AR-LSAT (Zhong et al., 2022).
+The logic-LLM approaches often combine Chain-of-thought and symbolic solvers in an agentic-LLM approach (Plaat et al., 2025), or transform the problem in an intermediate formal language.
+The use of LLMs for generating formal languages is further discussed in section 4.4.
+More recently, LLMs are also shown to struggle with analogies and analogical reasoning (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990), a topic that is related to self-reflection (see Sections 3.3.3 and 4.3.3).
+The question whether LLMs can reason prompted the development of the GSM8k benchmark (Cobbe et al., 2021) in 2021, a benchmark of some of easy reasoning problems (of the type: “Question: Natalia sold clips to 48 of her friends in April, and then she sold half as many clips in May.
+How many clips did Natalia sell altogether in April and May?
+Answer: 72”).
+This benchmark was specifically designed to test if LLMs can answer basic mathematical multi-step reasoning questions, or if they can not reason at all.
+Initially, performance was poor, but a year later Wei et al.
+(2022b) showed that by providing multi-step examples in the prompt—a chain of thought—much better performance was possible.
+Many works followed, and reasoning in language models has become an active area of study.
+In this survey, we focus on prompt-based multi-step reasoning algorithms.
+After starting with math word problems, work has more recently been extended to combinatorial games, puzzles, robotics, logic, and other fields of reasoning.
+At the end of this survey, we discuss connections to other fields, such as self-reflection and in-context reinforcement learning.
+Before we discuss the works on reasoning, this section reviews background terminology on LLMs.
 Our overview is brief.
-Excellent surveys on LLMs are, for example, Minaee et al.
+Excellent recent general surveys on LLMs are, for example, Minaee et al.
 (2024) and Zhao et al.
 (2023).
-We discuss the generic training pipeline for LLMs, we discuss how in-context learning works, and we discuss the reasoning pipeline.
-We start with the generic language model training pipeline.
-2.1 Training Pipeline Language Model
+Other works focus on the nature of reasoning and its definition (Yu et al., 2023; Chu et al., 2024a; Huang and Chang, 2023), on evaluating logical reasoning in LLMs (Mondorf and Plank, 2024; Liu et al., 2023a; Pan et al., 2023), on reinforcement learning in reasoning (Xu et al., 2025), on reasoning in language (Yu et al., 2023), or on Chain-of-thought itself (Chu et al., 2024b).
+The papers in this survey were selected as follows.
+We started by selecting papers on the ability to solve math word benchmarks (as a proxy for reasoning ability), that contained the search terms reasoning and large language model in their title or abstract, with a focus on papers that reference the Chain-of-thought paper.
+Although multi-step reasoning with LLMs initially was aimed at solving math world problems, it is now wider, including benchmarks and approaches for computer code, game play, puzzles, robot movement, and webpage navigation (see Table 2).
+We selected recent papers (two years prior to the writing of the survey) that show experimental results on selected benchmark datasets.
+We focus on prompt-based, in-context learning, methods based on Chain-of-thought, that are used in reasoning LLMs such as OpenAI o1 and o3 (Xu et al., 2025; Wu et al., 2024a; Huang et al., 2024).
+We also include papers that work with external algorithms, finetuning or supervised learning that have contributed to the approaches.
+We discuss the generic training pipeline for LLMs, we discuss how in-context learning works, and we discuss commonly used benchmarks. We start with the generic language model training pipeline.
+2.2. Language Model Training Pipeline
 LLMs are typically constructed in a sequence of stages, from data preparation, through training, to inference.
 The training pipeline for most LLMs is quite elaborate.
-We will now list a pipeline of the most common stages, based on the survey by Minaee et al.
+We will now list a brief pipeline of the most commonly used stages, based on the survey by Minaee et al.
 (2024).
-1.
-Acquire a large, general, unlabeled, high-quality text corpus. Some considerations on the selection of the texts are discussed in Brown et al. (2020).
-2.
-Pretrain the transformer model (Vaswani et al., 2017) on this large corpus.
-This step yields a generalist model.
-The pretraining is done using a self-supervised approach on the unlabeled dataset (text corpus).
-3.
-Finetune the general model to a specific (narrow) task.
-This can be done using supervised-learning with a new labeled dataset consisting of prompts and answers (supervised finetuning, SFT) (Wei et al., 2022a; Minaee et al., 2024), specific for the task at hand.
-(A small number of papers in this survey work in the finetuning stage.)
-4.
-Instruction tuning is a form of finetuning on a labeled dataset of instruction prompts and corresponding outputs, to improve instruction following, and thus the usefulness of models.
-5.
-Align the finetuned model with user expectations (preference alignment).
-The goal of this stage is to improve the model to give more ethically and socially acceptable answers.
-The machine learning method that is used in this stage can be, for example, Reinforcement Learning with Human Feedback (Ouyang et al., 2022) or Direct Preference Optimization (Rafailov et al., 2024).
-- 6.
-7.
-Inference & In-context learning can be used to train the model to provide the correct answers without changing parameters (Dong et al., 2022; Brown et al., 2020).
-By providing a prompt that contains a small number of examples together with a question, prompt learning is a form of few-shot learning.
-This is the stage in which most of the papers of this survey work, and that is familiar to all general users of ChatGPT.
-Most of the reasoning methods that we discuss in this survey work in stage 7: in-context learning, using prompts for the LLM to perform a complex multi-step reasoning task.
+In training an LLM, the first stage is to acquire a large, general, unlabeled, high-quality text corpus.
+Some considerations on the selection of the texts are discussed in Brown et al.
+(2020).
+The next stage is pretraining the transformer model (Vaswani et al., 2017) on this large corpus.
+This stage yields a generative language model.
+Pretraining is done using self-supervised autoregressive training on the unlabeled dataset (text corpus).
+Then the general model is finetuned to a specific (narrow) task.
+This can be done using supervised learning with a new labeled dataset consisting of prompts and answers (supervised finetuning, SFT, for example with low-rank optimization (Hu et al., 2022; Wei et al., 2022a; Minaee et al., 2024), or reinforcement learning (Sutton and Barto, 2018; Chu et al., 2025), specific for the task at hand.
+A small number of the surveyed papers also address the finetuning stage.
+A specific form of finetuning is instruction tuning, to improve instruction following for a certain task.
+Instruction tuning is supervised by a labeled dataset of instruction prompts and corresponding outputs.
+Depending on the purpose of the model the next step is alignment of the finetuned model with user expectations (preference alignment).
+Alignment is usually performed to ensure that the model produces more ethically and socially acceptable answers, preventing, for example, hate speech.
+The machine learning method that is commonly used in this stage Reinforcement Learning with Human Feedback (Ouyang et al., 2022) or Direct Preference Optimization (Rafailov et al., 2024).
+Optionally, model training can be optimized to improve cost-effectiveness, for example, mixed precision training (Micikevicius et al., 2018), quantization (Jacob et al., 2018), or knowledge distillation (Xu et al., 2024b; Gu et al., 2023).
+Once the model has been trained in the steps described above, it can be used further in the inference stage.
+Here, the model is used to provide an answer to the prompt.
+The inference-time stage is post-training, no model parameters are changed anymore (Dong et al., 2023; Brown et al., 2020); in-context learning, or prompt-learning, takes place in this stage.
+Due to its low barrier of entry, without the need for expensive training or finetuning, this form of advanced prompt engineering has become a popular training method.
+This is the stage on which most of the surveyed papers focus, using prompts for the LLM to perform a complex multi-step reasoning task.
 The following section provides a brief introduction to in-context learning.
-2.2 In-Context Learning
-In LLMs beyond hundreds of billions of parameters a new kind of learning has emerged, that is called in-context learning or prompt-learning (Brown et al., 2020).
-It occurs at inference time, and is often able to give good results with few examples; it is a form of few-shot learning.
-The large size of the model, containing rich and general knowledge is enabling this new type of few-shot learning (see Dong et al.
+2.3. In-Context Learning
+In large models, beyond hundreds of billions of parameters, a new kind of learning has emerged, that has been called in-context learning or prompt-learning (Brown et al., 2020).
+It occurs not when the model is trained, but when it is used, at inference time.
+Since no parameters are changed in this stage, it is not a model training stage; in-context-learning “learns” inside the context, or prompt, using information that is already encoded in the trained model parameters and the prompt, not by training the model anymore.
+In-context learning is often able to give good results with few examples, so-called few-shot learning, learning from the few examples in combination with the knowledge of the model.
+The large size of the model, containing rich and general knowledge, is enabling the few-shot learning (see Dong et al.
 (2023) for a survey).
-A prompt, consisting of a piece of demonstration context, is concatenated with a query question, and is given to the language model for prediction (Liu et al., 2023).
-For example, when the task is emotion recognition in a social media post, “I missed the bus today,” can be followed by “I felt so [___]”.
-Alternatively, for translation, we could follow “I missed the bus today,” by “French: [___]” (Liu et al., 2023).
+In in-context learning, a prompt, consisting of a piece of demonstration context, is concatenated with a query question, and is given to the language model, for text generation (Liu et al., 2023b).
+For example, when the task is emotion recognition in a social media post, “I missed the bus today,” can be followed by “I felt so [___]”, and the model could answer with “bad”.
+Alternatively, for translation, we could follow “I missed the bus today,” by “French: [___]” to request a translation (Liu et al., 2023b).
 The prompt contains background information that is recognized by the model, selecting the desired model context.
-In-context learning works when language models contain enough knowledge, allowing them to generalize on the examples provided in the prompt.
-Prompts that contain a few examples are said to perform few-shot learning. Prompts that contain only instructions without examples are said to perform zero-shot learning.
-In-context learning takes place at inference time, after the computationally intensive stages where parameters have been pretrained and finetuned, when the model is queried by the user to provide answers.
+In-context learning works when language models contain enough knowledge, allowing them to generalize on the (few) examples provided in the prompt.
+Contexts that contain a few examples are said to perform few-shot learning.
+Contexts that contain only instructions with zero examples are said to perform zero-shot learning.
+In-context learning takes place at inference time, after the computationally intensive training stages where parameters have been pretrained and finetuned, when the model is queried by the user to provide answers.
 No parameters are changed anymore with in-context learning.
 This is quite different from the common approach in supervised deep learning—or self-supervised deep learning—where large datasets are used during training to update model parameters with backward propagation in lengthy and costly training epochs (Goodfellow et al., 2016).
-Common approaches to few-shot learning, such as metalearning, do include training and finetuning of parameters to achieve generalization, and are computationally expensive (see, for example, Finn et al.
-(2017) or Huisman et al.
-(2021); Hospedales et al.
-(2021) for a survey).
+Indeed, in-context learning takes place fully at inference time, no parameters are trained, instead, learning now refers to adjusting the answers to the examples in the prompt and the internal knowledge acquired during training.
+Common approaches to few-shot learning, such as metalearning, do include training and finetuning of parameters to achieve generalization, and are computationally expensive (see, for example, (Finn et al., 2017) or (Huisman et al., 2021; Hospedales et al., 2021) for a survey).
+In-context learning, in comparison, is computationally cheap, and it has become a popular research approach..
 Prompts provide a user-friendly interface to LLMs.
-The success of in-context learning tends to be quite sensitive to the way a prompt is formulated; a new field called prompt engineering has emerged to optimize the usefulness of in-context learning by learning how to make them do what we want (Radford et al., 2019; Wei et al., 2022a; Giray, 2023; Sahoo et al., 2024).
-2.3 Reasoning Pipeline
-Reasoning problems are also solved with a pipeline of stages.
-A typical approach to solving a complex problem is to subdivide it into smaller steps and solve those.
-This approach is related to divide and conquer (Bellman, 1966).
-New steps are (1) generated, (2) evaluated, and the number of steps that are generated and searched is (3) controlled in some way.
-The in-context reasoning approaches that we survey follow a general three-stage pipeline (Madaan et al., 2023):
-1.
-Generate: generation of steps by the model,
-2.
-Evaluate: evaluation of the predicted steps by an evaluator,
-3.
-Control: control of the number of steps that are generated and how deep ahead the reasoning process will look.
-This three-stage pipeline will be the basis of our taxonomy. But first, we will look at benchmarks.
-3 Benchmarks
+The success of in-context learning tends to be quite sensitive to the way a prompt is formulated; a new field called prompt engineering has emerged to help human users learn how to make LLMs do what we want them to do (Radford et al., 2019; Wei et al., 2022a; Giray, 2023; Sahoo et al., 2024).
+The current survey thus discusses advanced prompt engineering methods.
+2.4. Multi-step Reasoning Benchmarks
 Progress in artificial intelligence is measured by benchmarks.
 Benchmarks define the goal that researchers aim to achieve in their experiments.
-In natural language processing, a wide array of benchmarks exists to measure progress, such as on question answering (for example, CommonsenseQA (Talmor et al., 2018)), word prediction (for example, LAMBADA (Paperno et al., 2016)), translation (for example, WMT’22 (Kocmi et al., 2022)), language understanding (for example, GLUE (Wang et al., 2018, 2019)), and text summarization (for example, Xsum (Narayan et al., 2018)).
-Transformer architectures were first popularized by encoder models such as BERT (Devlin et al., 2018), for named entity recognition and classification tasks.
-Subsequently, decoder models such as GPT 2-4 (Radford et al., 2019; Brown et al., 2020; Achiam et al., 2023) showed impressive progress on natural language benchmarks.
+In natural language processing, a wide array of benchmarks exist to measure progress, such as on question answering (for example, CommonsenseQA (Talmor et al., 2019)), word prediction (for example, LAMBADA (Paperno et al., 2016)), translation (for example, WMT’22 (Kocmi et al., 2022)), language understanding (for example, GLUE (Wang et al., 2018, 2019)), and text summarization (for example, Xsum (Narayan et al., 2018)).
 The field of LLMs is quite active.
-Many different benchmarks exist, and listing a comprehensive overview of all relevant benchmarks is beyond the scope of this survey.
-We will mention relevant benchmarks for testing the reasoning abilities of LLMs.
-Following Wei et al.
-(2022b), these are all math word problem benchmarks.
-The benchmark that is most frequently associated with reasoning by LLMs is a dataset of grade school math word problems GSM8K (Cobbe et al., 2021).
-GSM8K was created by humans, with an aim of high quality, high diversity, moderate difficulty, and solutions in natural language.
-Other benchmarks are the SVAMP varying structures benchmarks (Patel et al., 2021), the ASDiv dataset of diverse math problems (Miao et al., 2021), the AQuA dataset of algebraic word problems (Ling et al., 2017), and the MAWPS benchmark (Koncel-Kedziorski et al., 2016).
-We will now briefly discuss these benchmarks; the baseline performance that we quote is from Wei et al. (2022b).
-GSM8K
-To test reasoning skills, the Grade School Math problem dataset (GSM8K) was developed for testing LLMs (Cobbe et al., 2021).
+We will mention relevant benchmarks for testing the multi-step reasoning abilities of LLMs.
+The research on reasoning with LLMs started with math word problem benchmarks.
+The benchmark that is most frequently associated with multi-step reasoning with LLMs is the dataset of grade school math word problems GSM8K (Cobbe et al., 2021).
+GSM8K was created with the aim of providing high quality, high diversity, moderate difficulty, problems and solutions in natural language.
 It consists of 8500 human-written math problems.
-Language models struggled to achieve good performance on this dataset (pre Chain-of-thought).
-An example of a math word task is:
-Problem: Beth bakes 4, two dozen batches of cookies in a week. If these cookies are shared amongst 16 people equally, how many cookies does each person consume?
+Language models struggled to achieve good performance on this dataset before Chain-of-thought was introduced.
+An example of a math word task follows.
+Problem: Beth bakes 4 trays with two dozen batches of cookies in a week.
+If these cookies are shared amongst 16 people equally, how many cookies does each person consume?
 Answer: .
-The baseline performance of GPT-3 175B is 15.6% accuracy. In comparison, the performance of Chain-of-thought is 46.9% accuracy.
-ASDiv
-The Academia Sinica Diverse MWP Dataset (ASDiv) (Miao et al., 2021) is specifically designed for high diversity in problem types, formats and difficulty levels.
-It consists of 2305 problems.
-An example problem is:
-Problem: A sandwich is priced at . A cup of pudding is priced at . Tim bought 2 sandwiches and 4 cups of pudding. How much money should Tim pay?
-Answer: .
-The baseline performance of GPT-3 175B is 70.3% accuracy. The performance of Chain-of-thought is 71.3% accuracy.
-MAWPS
-The Math Word Problem Repository (MAWPS) (Koncel-Kedziorski et al., 2016) allows for the construction of datasets with particular characteristics by selecting different categories of problems.
-The dataset consists of 3320 problems.
-An example is:
-Problem: Rachel bought two coloring books. One had 23 pictures and the other had 32. After one week she had colored 44 of the pictures. How many pictures does she still have to color?
-Answer: .
-The baseline performance of GPT-3 175B is 72.7% accuracy. The performance of Chain-of-thought is 87.1% accuracy.
-SVAMP
-The Simple Variations on Arithmetic Math word Problems dataset (SVAMP) was designed by Patel et al.
-(2021).
-It consists of 1000 problems, curated from variations of ASDiv-a (Miao et al., 2021) and MAWPS (Koncel-Kedziorski et al., 2016).
-An example problem is:
-Problem: Jack had 8 pens and Mary had 5 pens. Jack gave 3 pens to Mary. How many pens does Jack have now?
-Answer: .
-The baseline performance of GPT-3 175B is 65.7% accuracy. In comparison, the performance of Chain-of-thought is 68.9% accuracy.
-AQuA
-The Algebraic Question Answering dataset (Ling et al., 2017) is a large dataset of 100,949 questions, answers, and rationales.
-The dataset is based on a combination of a smaller seed dataset and crowdsourcing.
-An example question is:
-Question: Two trains running in opposite directions cross a man standing on the platform in 27 seconds and 17 seconds respectively and they cross each other in 23 seconds.
-The ratio of their speeds is: Options: A) 3/7 B) 3/2 C) 3/88 D) 3/8 E) 2/2
-Answer: B.
-The baseline performance of GPT-3 175B is 24.8% accuracy. The performance of Chain-of-thought is 35.8% accuracy.
-There is a wide variety of benchmarks, and there is a wide variety of performance in benchmarks.
-Some are easily solvable by current LLMs, and some are significantly harder.
-Benchmark design is an important part of the field of reasoning in LLMs.
-Currently the GSM8K benchmark is popular; baseline model performance is weak, and reasoning prompts can substantially improve performance.
-As performance on GSM8K improves, different (harder) benchmarks will become popular.
-4 Selection of Papers
-The papers in this survey were selected as follows.
-Baseline LLMs have difficulty solving math word problems, specifically on benchmarks listed in the previous section.
-We take the ability to solve those benchmarks as a proxy for reasoning ability.
-We initially performed a literature search for papers that use these benchmarks, and that contain the search terms reasoning and large language model in their title or abstract.
-We also searched for papers that referenced the Chain-of-thought paper.
-The resulting papers were curated based on recency, relevance, substance, and novelty.
-We favor recent papers (two years prior to the writing of the survey), related to the Chain-of-thought approach of generating intermediate reasoning steps, that solve tasks such as math word problems, and that work by prompt-based in-context learning.
-We also include some papers that work by finetuning or supervised learning that relate to, or inspire, the Chain-of-thought approaches.
-Furthermore, we include approaches outside math word problems that showed interesting approaches to reasoning, such as applications in coding and autonomous agents, because of their approach to grounding.
-5 Prompt Generation, Evaluation and Control
-This survey examines how an architecture that is good at System 1 tasks can be prompted to solve System 2 tasks.
-The Chain-of-thought paper showed how a simple command could prompt an LLM to perform reasoning steps, yielding much better performance in math word problems.
-Since then much research has further explored this approach, trying to build the ultimate general problem solver for System 1 and System 2 problems.
-Following the pipeline of Section 2.3, the prompts must (1) generate the reasoning steps, (2) evaluate the answer to the steps, and (3) control the number of steps that are generated, the shape (or complexity) of the reasoning process must be controlled.
-We will now briefly discuss the three stages.
-Please refer to Figure 1 for a diagram of the different approaches for the generation, evaluation, and control of reasoning steps, and to Table 1.222We show the approaches in the Figure in their main category only.
+Other benchmarks of similar math word problems are the SVAMP varying structures benchmarks (Patel et al., 2021), the ASDiv dataset of diverse math problems (Miao et al., 2020), the AQuA dataset of algebraic word problems (Ling et al., 2017), and the MAWPS benchmark (Koncel-Kedziorski et al., 2016).
+Table 1 summarizes the accuracy of Chain-of-thought on these basic math word problems, against the baseline of GPT-3 175B as LLM (Wei et al., 2022b), as percentage of benchmark questions answered correctly.
+| Benchmark | GPT-3 175B | Chain-of-thought (Wei et al., 2022b) |
+| GSM8K | 15.6 | 46.9 |
+| ASDiv | 70.3 | 71.3 |
+| MAWPS | 72.7 | 87.1 |
+| SVAMP | 65.7 | 68.9 |
+| AQuA | 24.8 | 35.8 |
+We see that Chain-of-thought performs well against the baseline of GPT-3 on some benchmarks, but there is certainly room for further improvement on others.
+In addition to the initial set of math word benchmarks, further reasoning approaches have been introduced that test performance in other fields of reasoning.
+Benchmarks have been developed on Advanced mathematical questions (Codeforces, ), Computer code comprehension (Human evaluation, Spider (Yu et al., 2018), Transcoder (Roziere et al., 2020)), Robotic movement (Alfworld (Shridhar et al., 2021), Kitchen (Ahn et al., 2022)), Puzzle solving (Game24 (Yao et al., 2024)), Creative writing (Yao et al., 2024)), Gaming (Checkmate problems(Yang et al., 2024), MineCraft (Fan et al., 2022)), and Webpage navigation (WebShop (Yao et al., 2022)).
+These benchmarks are used by other approaches in our survey, as we will see in more detail in Section 3.
+The scope of this survey is limited to multi-step reasoning approaches.
+Other studies have published more challenging benchmarks, to study the performance of Chain-of-thought and self-reflection in LLMs in puzzles and (logic) games (Ruoss et al., 2024; Paglieri et al., 2024; Wen et al., 2025).
+It can be difficult to directly compare different benchmarking results, since there may be differences in the way the experiments were conducted.
+Kamoi et al.
+(2024) provide a critical analysis of benchmarking problems in LLM reasoning.
+Despite these challenges, general conclusions can be drawn.
+Section 4.1 provides guidance on how to choose reasoning approaches for different problem types.
+3. Step Generation, Evaluation and Control
+This survey examines how LLMs based on the transformer architecture can be prompted to solve multi-step reasoning tasks.
+The Chain-of-thought paper shows how a simple command could prompt an LLM to perform reasoning steps, yielding much better performance in math word problems.
+Since then much research has further explored this approach, trying to build stronger general problem solvers for other types of reasoning problems.
+A typical approach to solve a complex problem is to subdivide it into smaller steps and to solve those.
+This approach is related to classical divide and conquer (Bellman, 1966).
+It consists of three stages.
+New steps are (1) generated, (2) evaluated, and the search of the generated steps is (3) controlled in some way.
+The in-context reasoning approaches that we survey follow a general three-stage pipeline (Madaan et al., 2023):
+Generate: prompt the model for the generation of steps,
+Evaluate: prompt for evaluation of the generated steps,
+Control: prompt for control of the number of steps that are generated and how deep ahead the reasoning process will look.
+This three-stage pipeline is the basis of our taxonomy.
+We will now discuss the three stages more deeply; for ease of reference they are numbered according to the Subsection in which they are described in more detail (3.1, 3.2, 3.3).
+Please also refer to Figure 1, or Table 2, for a diagram of the categories and subcategories of different approaches for the generation, evaluation, and control of reasoning steps.111We show the approaches in the Figure in their main category only.
 Some approaches show innovations in two categories, and are shown twice.
-(Since all approaches have a generation, an evaluation, and a control aspect, all could in principle occur three times—all three columns can be found in Table 1).
-Prompt for Step Generation
-The first order of business is to create a prompt that instructs the LLM to generate reasoning steps.
+(Since all approaches have a generation, an evaluation, and a control aspect, all could in principle occur three times, and all three columns can be found in Table 2).
+(3.1) Generation
+The first stage is to create a prompt that instructs the LLM to generate reasoning steps.
 The problem must be split into substeps.
-This can be achieved with a problem-specific prompt that contains elements of the problem, such as: “First calculate how many marbles Mary had originally, then how many her friend had, and finally how many they had together.”
-In general, it is possible to prompt an LLM to fill in the blanks in a step-by-step fashion.
-In the papers that we will discuss, there are three main approaches for generating the step-by-step prompt.
-The prompt may be (1) handcrafted for the problem by the researchers (hand-written prompt), or (2) the prompt or prompts may come from an source that is external to the model, such as another model or a dataset (prompt using external knowledge), or (3) the model itself can be prompted to generate a (series of) prompt(s) to analyze the problem (model-generated prompt).
+This can be achieved with a problem-specific prompt that contains elements of the problem, such as: “First calculate how many marbles Mary had originally, then how many her friend had, and finally how many they had together.” In general, it is possible to prompt an LLM to fill in the blanks in a step-by-step fashion.
+In the papers that we discuss, there are three main approaches for generating the step-by-step prompt, numbered with the Subsection in which the approaches are described.
+First, the prompt may be handcrafted for the problem by the researchers: (3.1.1) hand-written prompt.
+Second, the prompt or prompts may come from a source that is external to the model, such as another model or dataset: (3.1.2) external knowledge-based prompt.
+Third, the model itself can be prompted to generate a (series of) prompt(s) to analyze the problem (3.1.3) model-generated prompt.
 As we will see, all three approaches have their advantages and disadvantages.
@@
 In some of the early papers (around 2022) it is the only stage of the pipeline that is performed.
-Prompt for Result Evaluation
-After the prompt has been generated and the model has answered it, the next step in the reasoning pipeline is to evaluate the answer.
+(3.2) Evaluation
+After the prompt has been generated and the model has answered it, the next stage is to evaluate the quality of this answer.
+Such evaluation is often necessary to improve and perform well on the benchmark.
 Again, we see three main approaches for substep evaluation.
-First, the steps may be evaluated by (1) the model itself (self-assessment).
-Second, (2) an external program can be used to evaluate the steps.
-For example, when the steps are expressed as computer code, an external interpreter or compiler can be used to check the validity and the outcome (tool-based evaluation).
-Finally, (3) an external model can be used, LLM or otherwise.
-For example, in robotics, an external physics model can determine if certain actions are physically possible (external model validation).
-Perform Control of Reasoning Steps
+First, the steps may be evaluated by the model itself: (3.2.1) self-assessment.
+Second, an external program can be used to evaluate the steps.
+For example, when the steps are expressed as computer code, an external interpreter or compiler can be used to check the validity and the outcome: (3.2.2) tool-based evaluation.
+Finally, an external model can be used, LLM or otherwise.
+For example, in robotics, an external physics model can determine if certain actions are physically possible: (3.2.3) external model validation.
+(3.3) Control
+The third stage is control.
 A reasoning process that consists of multiple steps is a sequential decision process (Littman, 1996).
@@
 Again, we distinguish three approaches.
-Most papers use (1) a greedy selection approach: a single prompt with a single chain of steps is generated, and these steps are directly executed and followed.
-The second approach (2) is to generate an ensemble strategy of reasoning steps, evaluate them, combine the individual results, and present them as the result of the ensemble.
-Finally, (3) a full tree-search or a reinforcement learning (RL) algorithm can be used as scaffolding.
-In this case, when a step is followed and evaluated, the LLM can roll back and try a different reasoning step.
-This is a breadth-first search approach (Plaat, 2020).
+Most papers use a (3.3.1) greedy selection approach: a single prompt with a single chain of steps is generated, and these steps are directly executed and followed.
+The second approach is to generate an (3.3.2) ensemble strategy of reasoning steps, evaluate them, combine the individual results, and present them as the result of the ensemble.
+Finally, a full tree-search or a (3.3.3) reinforcement learning (RL) algorithm can be used as scaffolding.
+In this case, when a step is followed and evaluated, the LLM can roll back and try a different reasoning step (Plaat, 2020).
 Going further, a full reinforcement learning approach can be used (Sutton and Barto, 2018; Plaat, 2022) to find an optimal policy for the sequential decision process.
-A full Markov Decision Process of state, action, transition, and reward function is specified, and step control can become a process where prompts are generated dynamically.
-| Approach | Domain | Step generation | Step evaluation | Step control |
-| Scratchpad (Nye et al., 2021) | math word | hand-wr/supervised | - | greedy/1prompt |
-| Chain-of-thought (Wei et al., 2022b) | math word | hand-written | - | greedy/1prompt |
-| ZS-CoT (Kojima et al., 2022) | math word | hand-written | - | greedy/1prompt |
-| Auto-CoT (Zhang et al., 2022) | math word | model-generated | - | clustering |
-| Complexity (Fu et al., 2022) | math word | hand-written | self-consistency | greedy/1prompt |
-| Self-ask (Press et al., 2022) | math word | external knowledge | LLM | multi-hop questions |
-| Self-verification (Weng et al., 2022) | math word | hand-written | back-verify | ensemble |
-| Self-consistency (Wang et al., 2022b) | math word | hand-written | majority | ensemble |
-| Codex (Chen et al., 2021) | code | - | tool-based | - |
-| Self-debugging (Chen et al., 2023) | code | hand-written | tool-based | greedy |
-| Fun-search (Romera-Paredes et al., 2024) | code | hand-written | tool-based | evolutionary algorithm |
-| LLaMEa (van Stein and Bäck, 2024) | code | hand-written | tool-based | evolutionary algorithm |
-| MathPrompter (Imani et al., 2023) | math | hand-written | tool-based | ensemble |
-| Program-of-thoughts (Chen et al., 2022) | math word | hand-written, Codex | Python+Consist. | decouple reason/compute |
-| Program-aided-language (Gao et al., 2023) | math word | hand-written, Codex | NLP/Python | ensemble |
-| Refiner (Paul et al., 2023) | math word | finetune | critic model | gen/crit feedback |
-| Self-corrector (Welleck et al., 2022) | math word | finetune | corrector model | gen/corr feedback |
-| Self-improvement (Huang et al., 2022a) | math word | finetune | self-assessment | CoT/consistency |
-| Say-can (Ahn et al., 2022) | robot | model-generated | external model | greedy |
-| Inner-monologue (Huang et al., 2022b) | robot | hand-written | various | greedy |
-| Self-taught-reasoner (Zelikman et al., 2022) | math word | finetune | augmentation | greedy/feedback |
-| Least-to-most (Zhou et al., 2022) | math word | hand-written | self-assessment | curriculum |
-| Progressive-hint (Zheng et al., 2023) | math word | model-generated | self-assessment | stable prompt |
-| Self-refine (Madaan et al., 2023) | math word | model-generated | self-assessment | greedy/feedback |
-| Tree-of-thoughts (Yao et al., 2024) | puzzles | model-generated | self-assessment | BFS/DFS |
-| Buffer-of-thoughts (Yang et al., 2024) | math word | thought template | self-assessment | buffer manager |
-| Beam-search (Xie et al., 2024) | math word | model-generated | self-assessment | Beam Search |
-| ReAct (Yao et al., 2022) | action | external knowledge | self-assessment | reinforcement learning |
-| Reflexion (Shinn et al., 2024) | decision | model-generated | ext model | reinforcement learning |
-| Voyager (Wang et al., 2023) | Minecraft | model-generated | Minecraft | reinforcement learning |
-Domain
-Many papers are applied to math word problems (natural language descriptions of math problems).
-Math problems were the original inspiration for the experiments with reasoning in LLMs.
-Other application domains include autonomous agents, robotic movement, generating computer programs, and playing computer games.
-We will discuss these in more detail with the individual approaches.
+In general a Markov Decision Process of state, action, transition, and reward function can be specified, and step control can become a process where prompts are generated dynamically under the control of an external RL algorithm, or as in-context reinforcement learning (ICRL) (Lee et al., 2023).
+Approach Domain (3.1) Step generation (3.2) Step evaluation (3.3) Step control Result Scratchpad (Nye et al., 2022) math word hand-wr/superv - greedy/1prompt PolyEval +19%, Python +21% Chain-of-thought (Wei et al., 2022b) math word hand-written - greedy/1prompt GSM8K +39%, SVAMP +10%, ASDiv +2%, AQuA +11%, MAWPS +14%, CSQA +1.8%, StrategyQA +0.2% ZS-CoT (Kojima et al., 2022) math word hand-written - greedy/1prompt MultiArith =89%, GSM8K =70% Auto-CoT (Zhang et al., 2023) math word model-generated - clustering MultiArith +0.3%, GSM8K +1%, AddSub +3.5%, AQuA +0.7%, SingleEq +0.4%, SVAMP +0.6%, CSQA +1%, StrategyQA +0%, Letter +0.7%, Coin +2.7% Complexity (Fu et al., 2022) math word hand-written self-consistency greedy/1prompt GSM8K +7%, MultiArith +3%, Penguins +3% Self-ask (Press et al., 2023) math word ext knowledge LLM multi-hop q.
+Bamboogle =60%, 2Wiki =40%, Musique =15% Self-verification (Weng et al., 2023) math word hand-written back-verify ensemble GSM8K +4%, SingleEq +2%, AddSub +4%, MultiArith +3%, AQuA +3%, SVAMP +1% Self-consistency (Wang et al., 2023b) math word hand-written majority ensemble GSM8K +18%, SVAMP +11%, AQuA +12%, StrategyQA +6%, ARC-c +4% Codex (Chen et al., 2021) code - tool-based - HumanEval =70% Self-debugging (Chen et al., 2024) code hand-written tool-based greedy Spider +9%, MBPP +12%, TransCoder +12% Fun-search (Romera-Paredes et al., 2024) code hand-written tool-based evolutionary alg cap set 8 =512 LLaMEa (van Stein and Bäck, 2024) code hand-written tool-based evolutionary alg BBOB +11% MathPrompter (Imani et al., 2023) math hand-written tool-based ensemble MultiArith =92% Program-of-thoughts (Chen et al., 2023) math word hand-wr, Codex Python+Consist.
+split reason/cmput GSM8K =71%, SVAMP =85%, ASDIV =85%, AddSub =92%, MultiArith = 99% Program-aided-lang (Gao et al., 2023) math word hand-wr, Codex NLP/Python ensemble GSM8K =72%, SVAMP =79%, ASDIV =79%, SingleEQ =96%, SingleOP =94%, AddSub = 92%, MultiArith = 99%, Penguins = 93% Refiner (Paul et al., 2024a) math word finetune critic model gen/crit feedback SVAMP =72%, GSM8K =78% Self-correction (Welleck et al., 2023) math word finetune corrector model gen/corr feedback MathProgSynth =24%, LexConstrGen =98%, ToxicityControl =0.0% Self-improvement (Huang et al., 2023) math word finetune self-assessment CoT/consistency GSM8K =82%, DROP =83%, ARC-c =90%, OpenBookQA =94%, ANLI-A3 =68% Say-can (Ahn et al., 2022) robot model-generated external model greedy Kitchen =31% Inner-monologue (Huang et al., 2022) robot hand-written various greedy TableTop =90%, Kitchen =60% Self-taught-reasoner (Zelikman et al., 2022) math word finetune augmentation greedy/feedback CommonsenseQA =72% Least-to-most (Zhou et al., 2023) math word hand-written self-assessment curriculum SCAN =99% Progressive-hint (Zheng et al., 2023) math word model-generated self-assessment stable prompt AddSub +2%, MultiArith +0%, SingleEQ +2%, SVAMP +3%, GSM8K +5%, AQuA +1% Self-refine (Madaan et al., 2023) math word model-generated self-assessment greedy/feedback Sentiment +32%, Dialog +49%, CodeOptim +8%, CodeRead +28%, MathReason +0%, AcronymGen +25%, ConstrainedGen +30% Tree-of-thoughts (Yao et al., 2024) puzzles model-generated self-assessment BFS/DFS Game24 =74%, CreativeWriting , Crossword Buffer-of-thoughts (Yang et al., 2024) math word thought template self-assessment buffer manager Game24 +11%, GeoShapes +20%, Checkmate +51% Algorithm-of-thoughts (Sel et al., 2023) puzzles model-generated self-assessment in-context RL GSM8K =89%, StrategyQA =84%, Crossword Beam-search (Xie et al., 2024) math word model-generated self-assessment Beam Search GSM8K +6%, AQuA +9%, StrategyQA +5% ReAct (Yao et al., 2023) action external knowledge self-assessment reinf learning ALFWorld =34%, WebShop =10% Reflexion (Shinn et al., 2024) decision model-generated ext model reinf learning HumanEval =91% Voyager (Wang et al., 2024a) Minecraft model-generated Minecraft reinf learning 15 x faster
 Taxonomy Table
-Table 1 lists the papers of this survey.
-They are listed by the domain they work on, the type of prompt generation, the evaluation of the result, and the control method.
-The approaches in the table are grouped, divided by horizontal lines.
-The first group, from Scratchpad to Self-ask, focuses on creating a prompt that generates the reasoning steps.
+Table 2 lists the main papers of this survey.
+We show the domain they work on, the type of prompt generation, the evaluation of the result, and the control method.
+These three categories of approaches—indicated by their Sections (3.1) generation, (3.2) evaluation, (3.3) control—are shown in the table as groups divided by horizontal lines.
+The first group in the Table, from Scratchpad to Self-ask, focuses on creating a prompt that generates the reasoning steps.
 The entries in the cells of this column are shown in bold, highlighting the focus of the approaches.
-The approaches in this group can be considered to be the start of the field of LLM-reasoning.
+The approaches in this group are the start of the field of LLM-reasoning.
 The Chain-of-thought approach is especially an inspiration for many works.
-The prompts are often written “manually” by the researchers, the steps are encoded in one prompt, and step control is greedy.
+The prompts are often written manually by the researchers for each problem; the steps are encoded in one prompt, and step control is greedy.
 There is no specific evaluation of the steps, other than comparing results to the benchmark.
-The Scratchpad approach is special in that it uses supervised learning, not prompt-learning; the work showed that LLMs can be made to generate internal reasoning steps by supervised learning, paving the way for the later prompt-based papers.
+The Scratchpad approach is special in that it uses supervised learning, not prompt-learning; the work showed that LLMs can generate internal reasoning steps by supervised learning, paving the way for in-context works.
 The second group, from Self-verification to Self-taught-reasoner, focuses on evaluation of the reasoning steps in the prompt.
 This column is shown in bold in the table.
-The approaches in this group aim to improve the Chain-of-thought results by reducing the error accumulation that occurs when multiple steps are taken in a reasoning chain.
+The approaches in this group aim to improve the Chain-of-thought results by reducing error accumulation that occurs when multiple steps are taken in a reasoning chain.
 A variety of step control methods is used by these approaches, which is discussed in more detail later.
-Note that not all approaches use natural language problems (often math word problems).
+Note that not all approaches use natural language problems.
 For example, the subgroup of Codex to Program-aided-language focuses on formal languages.
 They generate code or math equations, typically in Python, to formalize the steps of the reasoning problem, or as result of the task.
-LLMs are quite good at code generation, and these approaches typically achieve good performance.
+LLMs are quite good at code generation (Chen et al., 2021), and these approaches typically achieve good performance.
 The use of code also allows the approaches to call external programs such as interpreters and debuggers to evaluate the correctness of the reasoning steps that are generated.
-There is also a special subgroup, Refiner to Self-improvement, that does not use prompt learning but finetuning.
+There is also a special subgroup, Refiner to Self-improvement, that uses finetuning in addition to prompt learning.
 Here, new data is generated based on reasoning exemplars, which is then used to further train the model.
@@
 The laws of physics are learned and used to ground the reasoning steps in reality (to reduce hallucination).
-The third group, Least-to-most to Voyager, addresses step control (approaches shown in bold in this column).
+The third group, Least-to-most to Voyager, addresses step control (shown in bold in this column).
 Whereas in the previous approaches the reasoning steps are written in a single, static, prompt, these approaches generate the steps in multiple, dynamic, prompts.
@@
 The control methods range from simple greedy and depth-first search to elaborate beam search and reinforcement learning schemes.
-In summary, we see a diverse array of methods that often achieve high performance in reasoning about their respective domains.
-To better understand the approaches, let us discuss the techniques in more detail, starting with the generation of steps.
-5.1 Generation of Steps
-Originally, LLMs performed poorly on math word problems (GSM8K (Cobbe et al., 2021)).
-Some different approaches were tried, for example scaling up the size of the LLM (Rae et al., 2021).
+The last column of the Table summarizes reported benchmark results.
+The ’=’ symbol indicates absolute scores on the benchmarks, while ’+’ indicates relative improvement in percentage points over a baseline LLM, typically GPT-3.5.
+Results vary strongly, both between approaches and within a single approach between benchmarks.
+Also, different LLMs were used, from early stage to more mature models, open and commercial, and the baselines differ.
+For some benchmarks, such as the Creative Writing benchmark in Tree-of-thoughts, results are best reported qualitatively.
+The source papers provide more measurement details, Section 4.1 discusses when to which approach to use for different applications.
+In conclusion, we see a diverse array of methods that often achieve high performance in reasoning on their respective domains.
+To better understand the approaches, we discuss them in more detail, starting with the generation of steps.
+3.1. Generation of Steps
+Originally, LLMs performed poorly on math word problems such as GSM8K (Cobbe et al., 2021).
+Different approaches were tried unsuccessfully, for example scaling up the size of the LLM (Rae et al., 2021).
 The LLM architecture, based on transformers, is designed to produce a single token.
 When we prompt such an architecture to produce an answer, it does so.
-What we should do is prompt it to follow intermediate steps, answer those, and thus work towards the final answer, just as a student is taught to break down a complex problem into smaller steps.
-We should take the model by its hand and teach it to write down the intermediate steps, and combine the intermediate results (Nye et al., 2021).
+What we should do instead, is to prompt it to follow intermediate steps, answer those, and thus work towards the final answer, just as a student is taught to break down a complex problem into smaller steps.
+We should guide the model to explicitly produce intermediate steps, and combine the intermediate results.
 This idea was used by Nye et al.
@@
 Figure 2 shows an example.
-On experiments with addition, polynomial evaluation, and Python code execution, versions that produced the intermediate steps on a scratchpad performed considerably better than versions that did not.
-If supervised learning can produce intermediate steps, would prompt-learning be able to do so too?
-5.1.1 Hand-written Prompt
+On experiments with addition, polynomial evaluation, and Python code execution, versions that produced the intermediate steps on a scratchpad performed considerably better than versions that did not, going from 35% to 95%, from 32% to 51%, and from 30% to 42% accuracy, respectively.
+If supervised learning can produce intermediate steps, would prompt learning be able to do so too?
+3.1.1. Hand-written Prompt
 This question was studied by Wei et al.
@@
 A basic way to instruct an LLM to generate steps by prompt-learning is to manually write a prompt for the large language model to follow the reasoning steps.
-They showed in their Chain-of-thought paper that with such a prompt the LLM follows such intermediate steps.
-When the LLM is prompted to rephrase information from the question as intermediate reasoning steps in its answer, the LLM performed much better than when it was prompted to answer a math problem directly, without reproducing the information from the question in its answer.
-The example from the Chain-of-thought paper is shown in Figure 3 Wei et al.
-(2022b).
-Performance figures were given in Section 3 on benchmarks.
-The substantial performance improvement by Chain-of-thought has caused much excitement and has opened up further research on reasoning with LLMs.
-In the original Chain-of-thought paper the prompts were handwritten by the researchers for the individual types of problems, and evaluations are conducted with five different benchmarks (not by an LLM).333The Chain-of-thought idea is about prompt generation, not about the evaluation or the search control of the reasoning steps.
+When the LLM is prompted to rephrase information from the question as intermediate reasoning steps in its answer, the LLM performed much better than when it was prompted to answer a math problem directly, without reproducing information from the question in its answer in multiple steps.
+The example from the Chain-of-thought paper is shown in Figure 3.
+Table 1 shows that the largest accuracy increase is on GSM8K, from 16% to 47%.
+The idea that an LLM can be made to follow step-by-step instructions, and the performance improvement by Chain-of-thought have caused much excitement and have opened up further research on reasoning with LLMs.
+In the original paper the prompts were handwritten by the researchers for the individual types of problems, and evaluations are conducted with benchmarks (not by an LLM).222The Chain-of-thought idea is about prompt generation, not about the evaluation or the search control of the reasoning steps.
 Hence, in Table 2 Chain-of-thought is labeled as greedy without an evaluation.
-In a later work the prompts were generated automatically by the LLM (Zhang et al., 2022).
+In a later work the prompts were generated automatically by the LLM (Zhang et al., 2023), and evaluated.
 Kojima et al.
 (2022) go a step further.
-They show that the simple addition of a single text to the prompt (Let’s think step by step) significantly improves performance.
-Since this text does not contain problem-related elements, this can be considered as a form of zero-shot learning.
-Figure 4 compares the approaches.
-Experiments further show that with this addition to the prompt, significant performance gains are achieved on a diverse set of reasoning benchmarks, including arithmetic, symbolic, and logical reasoning.
-5.1.2 Prompt using External Knowledge
-Chain-of-thought shows that an LLM gives better answers to complex problems when it is guided to take individual steps. Prompts are written manually, from scratch, by the researchers.
-We can use external information about the problem to improve the prompt.
+They show that the addition of a single text to the prompt (Let’s think step by step) significantly improves performance.
+Since this text does not contain problem-related elements, it is as a form of zero-shot learning.
+Figure 3 (third column) compares the approaches.
+Experiments further show that with this addition to the prompt significant performance gains are also achieved on other reasoning benchmarks, including arithmetic, symbolic, and logical reasoning (achieving 70% accuracy on GSM8K/PaLM when Self-consistency is also included).
+3.1.2. Prompt using External Knowledge
+Chain-of-thought prompts are written manually, by the researchers, an approach that does not scale.
+We can also use external information about the problem to improve the prompt.
 Press et al.
@@
 This difference is called the compositionality gap.
-They find that in GPT-3, as model size increases, the compositionality gap does not decrease: the single-hop question-answering performance improves faster than the multi-hop performance.
-This shows that while more powerful models memorize and recall more factual knowledge, no improvement in their ability to perform compositional reasoning occurs.
-They find that the ability to reason does not depend on the size of the model.
-Subsequently, a method called Self-ask is proposed, that asks elicitive follow-up questions (like Chain-of-thought, but with the follow up: prompt), see Figure 5.
-The model is then used to answer these follow-up questions.
+They find that in GPT-3, as model size increases, the single-hop question-answering performance improves faster than the multi-hop performance: while more powerful models memorize and recall more factual knowledge, no improvement in their compositional reasoning occurs.
+The ability to reason does not depend on the size of the model.
+Subsequently, a method called Self-ask is proposed, that asks elicitive follow-up questions (like Chain-of-thought, but with the follow up: prompt), that the model then answers.
 Self-ask can also use an external search engine to answer intermediate prompts, instead of the model.
-The model takes as input a compositional question which it decomposes.
 The initial subquestion is fed into the search engine, and the answer is processed by the model, which generates another subquestion, and so on, until it produces the final answer.
-The approach performs a few percentage points better than vanilla Chain-of-thought on three benchmarks that were specifically designed for multi-hop questions.
-5.1.3 Model-Generated Prompt
-In addition to manually writing prompts or using external information, we can also try to let the LLM itself study the problem to write the best reasoning-prompt, a form of self-improvement.
-An example of this approach is Auto-chain-of-thought (Zhang et al., 2022).
+Self-ask was tested on three benchmarks that were specifically designed for multi-hop questions.
+Although it performs only a few percentage points better than vanilla Chain-of-thought, it showed how external knowledge can be used in a reasoning setting.
+3.1.3. Model-Generated Prompt
+In addition to manually writing prompts or using external information, we can also let the LLM itself study the problem to write the best reasoning-prompt.
+An example of such self-improvement is Auto-chain-of-thought (Zhang et al., 2023).
 This approach builds on the observation by Kojima et al.
@@
 The constructed demonstrations augment the in-context prompt.
-The automatically generated prompts are reported to perform as well or better than the hand-written Chain-of-thought prompts on ten benchmarks using GPT-3.
+This approach also performed a few percentage points better than hand-written Chain-of-thought prompts, on ten benchmarks, using GPT-3 (see Table 2).
 Fu et al.
 (2022) introduce Complexity-based prompting.
-Inspired by Chain-of-thought and Self-consistency, this work studies which prompts achieve the best results on math word and other reasoning problems.
-Their work specifically studies the impact of the complexity of the reasoning chain, and introduces a related reasoning approach (Complexity-based prompting).
-They find that prompts with the largest complexity (the most reasoning steps) perform best.
-Further, they find that outputs (answers) with the highest complexity are the best.
-Complexity-based prompting achieves high performance on three math reasoning benchmarks.
-Another approach that uses model-generated prompts is Buffer-of-thoughts. We will discuss this approach in Section 5.3.3.
-5.2 Evaluation of Steps
-After discussing prompts for the generation of reasoning steps, the next stage in the reasoning pipeline (Section 2.3) is evaluation of the results of the steps, to reduce the error of multi-step reasoning chains.
+Inspired by Chain-of-thought and Self-consistency, their work specifically studies the impact of the complexity of the reasoning chain (the number of steps), and introduces a related reasoning approach (Complexity-based prompting).
+They find that prompts with the largest complexity perform best, and also that answers with the highest complexity are the best.
+Complexity-based prompting achieves somewhat higher performance on three math reasoning benchmarks: GSM8K improves 7 points, MathQA 6 points, and the Penguins benchmark from Big Bench Hard improve 3 percentage points.
+We see that the initial approaches showed larger improvements than the later approaches. It is time to look at another category of approaches, that focus on the evaluation of reasoning steps.
+3.2. Evaluation of Steps
+After discussing prompts for the generation of reasoning steps, the next stage in the generation/evaluation/control pipeline is evaluation of the results of the steps.
+This stage focuses on reducing error accumulation of multi-step reasoning chains.
 We will start with approaches where the same model performs step-generation and step-evaluation.
-5.2.1 Self-Assessment
+3.2.1. Self-Assessment
 When LLMs are prompted to perform reasoning steps, they perform a sequence of steps and predict multiple tokens.
-Performing a sequence of steps makes them sensitive to mistakes and vulnerable to error accumulation (Weng et al., 2022; Xiao et al., 2023a).
+Performing a sequence of steps makes them sensitive to mistakes and vulnerable to error accumulation (logical, factual, ethical, or otherwise) (Weng et al., 2023; Xiao et al., 2023a).
 Several methods have been developed to prevent error accumulation.
@@
 They note that human reasoning also suffers from the problem of accumulating errors, and that in human reasoning we frequently revisit our thought process to verify the accuracy of our reasoning steps.
-Thus, they propose to apply such a backwards self-verification approach.
-The LLM is prompted to use the conclusion of the Chain-of-thought reasoning chain as a condition for solving the original problem and then compare the answer going back to the original question.
+The LLM is prompted to use the conclusion of the Chain-of-thought reasoning chain as a condition for solving the original problem and then compare the answer, going back to the original question.
 The LLM is given variations of its own conclusion and is instructed to choose the one with the highest similarity to the original question.
@@
 (2024).) Experiments are reported on GPT-3 (Chen et al., 2021) and on Instruct-GPT (Ouyang et al., 2022).
-The performance of Chain-of-thought was improved by a few percentage points on arithmetic and general reasoning tasks.
-A popular related approach is called Self-consistency (Wang et al., 2022b).
-Self-consistency is a straightforward ensemble approach.
+The accuracy of Chain-of-thought was improved by a few percentage points on arithmetic and general reasoning tasks (GSM8K 65%, AQuA 48%, SVAMP 77%).
+A popular related approach is Self-consistency (Wang et al., 2023b).
+Self-consistency is a straightforward ensemble approach (a well-known machine learning technique to make a strong learner out of multiple weaker learners (Sagi and Rokach, 2018; Breiman, 2001)).
 Greedy single-path decoding is replaced by sampling diverse reasoning paths, evaluating them, and selecting the most consistent answer.
-Self-consistency asks the LLM to simply perform the same query multiple times, and takes the majority-vote of the answers.
+Self-consistency asks the LLM to simply perform the same query multiple times, and takes the majority-vote of the answers, or decoding paths.
 Self-consistency works since complex reasoning problems typically allow different reasoning paths that lead to the correct answer.
 Figure 4 summarizes the approach.
-Self-consistency has been evaluated on arithmetic reasoning, commonsense reasoning and symbolic reasoning, on a variety of LLMs, including GPT-3 (Tay et al., 2022; Brown et al., 2020; Thoppilan et al., 2022; Chowdhery et al., 2023).
-Self-consistency improves the performance of Chain-of-thought typically by 10-20 percentage points, and has been used as a baseline in many of the other approaches in this survey.
-(Self-verification also reports that performance is improved when used in combination with Self-consistency (Wang et al., 2022b) and with Program-aided-language (Gao et al., 2023).)
-5.2.2 Tool-based Validation
-Another possibility to improve the accuracy of evaluating the reasoning steps is to switch from a natural to a formal language.
+Self-consistency has been evaluated on arithmetic reasoning, commonsense reasoning and symbolic reasoning, on a variety of LLMs, including GPT-3 (Tay et al., 2023; Brown et al., 2020; Thoppilan et al., 2022; Chowdhery et al., 2023).
+Self-consistency further improves the performance of Chain-of-thought by 10-20 percentage points, and has been used as a baseline in many of the other approaches in this survey.
+3.2.2. Tool-based Validation
+Another approach to improve the accuracy of evaluating the reasoning steps is to switch from a natural to a formal language.
 The advantage of a formal language is that it is less ambiguous than a natural language.
@@
 Codex is able to generate correct programs from descriptions in natural language, such as commentary strings.
-Figure 7 shows examples that are produced by Codex.
 The work on Codex is used as a basis for further research on reasoning in LLMs.
 Human programmers, when writing code, typically follow a cycle of writing some code, executing it to look for errors, and then using the feedback to improve the code.
-This same approach is followed in the Self-debugging work (Chen et al., 2023).
-Self-debugging teaches a large language model to debug its generated program code via few-shot demonstrations.
-It follows the same steps of (1) code generation, (2) code execution, and (3) code explanation (see Figure 8).
-Self-debugging is able, without human feedback on the code’s correctness or error messages, to identify mistakes in the code that was generated by itself from investigating the execution results.
-Self-debugging can also provide an explanation of the generated code in natural language.
-It achieves strong performance on text-to-SQL generation, C++-to-Python transcoding, and text-to-Python generation.
-Several works use self-debugging to generate working code tuned for solving specific problems automatically, without human feedback.
+This step-by-step approach is followed in Self-debugging (Chen et al., 2024).
+It follows the same steps of code generation, code execution, and code explanation.
+Self-debugging is able to identify mistakes in its own code by investigating the execution results, and can also provide an explanation of the generated code, in natural language.
+It achieves strong performance: the text-to-SQL Spider benchmark improves 9 points, and the C++ to Python Transcoder benchmark improves by 12 percentage points.
+Several works generate working code tuned for solving specific problems automatically, without human feedback.
 Romera-Paredes et al.
 (2024) introduced FunSearch, an approach that integrates formal methods and LLMs to enhance mathematical reasoning and code generation.
-FunSearch is capable of producing functionally correct programs that adhere to specified requirements.
-It uses a genetic algorithm approach with multiple populations of candidate solutions (programs), which are automatically evaluated (using tools depending on the problem specification).
-In addition to the problem specification in the form of an evaluate function, also an initial program is given to the LLM in the first prompt.
-After evaluating a number of generated programs from the starting prompt, a new prompt using ‘best-shot prompting’ is created in an iterative fashion, combining a selection of sampled programs in a sorted list (ascending according to their evaluation score), and the LLM is requested to generate program .
+FunSearch uses a genetic approach with multiple populations of candidate solutions (programs), which are evaluated using a function depending on the problem specification.
+In addition to the evaluation function, also an initial program is given to the LLM in the first prompt.
+After evaluating a number of generated programs from the starting prompt, a new prompt is created, in an iterative fashion, combining a selection of sampled programs sorted according to their evaluation score, and the LLM is requested to generate a new program.
 Another work leverages evolutionary computation methods to generate and optimize evolutionary algorithms (van Stein and Bäck, 2024).
@@
 The approach uses LLMs to generate initial algorithmic structures, which are then refined through mutation and selection.
-This enhances the efficiency of algorithm design, particularly in fields requiring innovative and adaptive solutions.
+This enhances the efficiency of algorithm design, particularly in fields requiring innovative and adaptive solutions, improving accuracy on the Black-Box Optimization Benchmark (Hansen and Ros, 2010) (BBOB) suite by 11 pecentage points.
 A key difference between FunSearch and LLaMEA is that LLaMEA uses a sample-efficient elitism strategy by iteratively improving the best-so-far solution, requiring significantly fewer prompt evaluations than the large-population strategy proposed in FunSearch.
+Evolutionary approaches let the LLM discover new algorithms, solving existing problems in new ways, or solving entirely new problems.
+Another method, Evolution-of-heuristics (Liu et al., 2024), was proposed for evolving code snippets for guided local search to solve combinatorial optimization problems, such as the Traveling Salesperson Problem.
 To improve prompt-based reasoning, Codex is used in an ensemble approach named MathPrompter (Imani et al., 2023).
@@
 The results are compared, just like in Self-consistency and Self-verification, raising the confidence level in the results.
-MathPrompter achieved state-of-the-art results on the MultiArith dataset (78.7% 92.5%), evaluated on GPT-3 175B.
-Two other approaches that use a formal language are Program-of-thought (PoT) (Chen et al., 2022) and Program-aided-language (PAL) (Gao et al., 2023).
-Both approaches use the LLM to generate Python and then use a Python interpreter to evaluate the result.
-PoT and PAL are similar approaches.
-PoT uses benchmark-specific prompts; PAL uses generic prompts, and has been tested on more benchmarks and has been used in other approaches.
-Figure 9 illustrates the PAL approach.
+MathPrompter achieved state-of-the-art accuracy on the MultiArith dataset (78.7% 92.5%), evaluated on GPT-3 175B.
+Two other approaches that use a formal language are Program-of-thought (Chen et al., 2023) and Program-aided-language (Gao et al., 2023).
+Both approaches use the LLM to generate Python and then use an interpreter to evaluate the result.
+The approaches are similar although Program-aided-language uses generic prompts, and has been tested on more benchmarks.
+Figure 3 (fourth column) illustrates the Program-aided-language approach.
 When the evaluation of the reasoning steps is offloaded to the Python interpreter, decomposing the natural language problem into executable code-steps remains the only task for the LLM.
-(Earlier work in mathematical word problems, such as Ling et al.
-(2017), showed how to decompose a problem and reach an answer.) Gao et al.
+(Earlier work in math word problems showed how to decompose a problem and reach an answer (Ling et al., 2017).) Gao et al.
 (2023) provide extensive experimental evidence about the synergy between the neural LLM and the symbolic interpreter.
-Experiments are performed over 13 mathematical, symbolic, and algorithmic reasoning tasks, achieving more accurate results than much larger models.
-5.2.3 External Model Validation
-We have seen many successful examples of prompt-based in-context reasoning and evaluation.
-We will now look at related reasoning approaches that follow a more traditional parameter learning approach.
-We describe three natural language approaches that follow this route.
+Experiments are performed on 13 mathematical, symbolic, and algorithmic reasoning tasks, achieving more accurate results than much larger models (for example, Program-aided-language reported 72% on GSM8K and 93% on Penguins).
+3.2.3. External Model Validation
+We have seen many examples of succesful prompt-based in-context reasoning and evaluation (at inference time—where no parameters were changed).
+We will now look at reasoning approaches that follow a more traditional parameter training approach.
 All approaches evaluate the output of the model and generate corrective data.
@@
 Finetuning
-The Refiner approach (Paul et al., 2023) uses a generator model and a critic model to provide fine-grained feedback on reasoning errors.
-The generator generates multiple reasoning hypotheses, the critic evaluates results by randomly selecting a hypothesis for feedback.
-The generator model is finetuned based on its reasoning errors.
+The Refiner approach (Paul et al., 2024a) uses a generator model and a critic model that provide fine-grained feedback on reasoning errors.
+The generator generates multiple reasoning hypotheses, and the critic evaluates results by randomly selecting a hypothesis for feedback.
+The generator model is then finetuned based on its reasoning errors.
 A small supervised model is used to overcome the cold-start problem.
-Figure 10 shows an example of how the critic provides feedback to the generator.
-The approach is reported to work well on math word problems and synthetic natural language reasoning.
+The approach achieves 78% accuracy on GSM8K and also works well on related problems.
 Welleck et al.
-(2022) follow a similar approach in their Self-correction approach.
-The corrector is a separate model specialized in refining the outputs of the generator.
-Unlike Refiner, where the generator is finetuned based on the critic feedback, Self-correction finetunes the corrector to rectify errors in the hypotheses produced by the generator.
-A third finetuning approach is Self-improvement, by Huang et al.
-(2022a).
-Here too the base model data is augmented by LLM-generated rationales, and then finetuned.
-Noteworthy in all three finetuning approaches is that LLMs are capable of improving themselves by training on their own generated output, and that stability problems inherent in feedback loops are overcome.
+(2023) follow a similar approach which they call Self-correction.
+Here, the corrector is a separate model specialized in refining the outputs of the generator.
+Unlike Refiner, where the generator is finetuned based on the critic, Self-correction finetunes the corrector to rectify errors in the hypotheses produced by the generator.
+Self-corrector is not applied to math word problems, but to program synthesis, where a small corrector reduces toxicity to 0%.
+A third finetuning approach is Self-improvement (Huang et al., 2023).
+Here, too, the base model data is augmented by LLM-generated rationales, and then finetuned.
+They achieve 82% accuracy on GSM8K and similarly high scores on question answering and adversarial benchmarks.
+Noteworthy in all three finetuning approaches is that LLMs are capable of improving themselves by training on their own generated output, and that stability problems from feedback loops are overcome.
 Dataset Augmentation
 The final finetuning approach that we discuss uses dataset augmentation.
-An explicit intermediate reasoning is called a rationale.
-Rationale generation has been shown to be valuable for LLMs across diverse tasks such as mathematical and commonsense reasoning, code evaluation, social bias inference, and natural language inference (Zelikman et al., 2022).
-Zelikman et al.
-(2022) describe how reasoning steps are used to create rationales, that are then used to augment the dataset on which the model is finetuned.
-The approach is called Self-taught-reasoner.
-Figure 11 illustrates the approach.
-In Self-taught-reasoner, an augmentation dataset is created by attempting to solve the original dataset using the current model’s rationale generation ability in each iteration.
-Next, the dataset is augmented using rationalizations, using ground-truth answers to problems the model failed to solve.
-Finally, the large language model is finetuned on the combined dataset.
+In Self-taught-reasoner (Zelikman et al., 2022), an intermediate reasoning is generated, called a rationale.
+Rationales are shown to be valuable across diverse tasks such as mathematical and commonsense reasoning, code evaluation, social bias inference, and natural language inference.
+First an augmentation dataset is created by attempting to solve the original dataset.
+Next, the dataset is augmented using rationalizations and ground-truth answers to problems the model failed to solve.
+Finally, the model is finetuned on the combined dataset.
+Self-taught-reasoner performs comparably (72%) to finetuning a 30 times larger model on CommonsenseQA.
 Reasoning about Robot Behavior
-In addition to math word problems, prompt-based reasoning has also been used to reason about robot behavior.
+In addition to math word problems, computer code, and common sense, prompt-based reasoning has also been used to improve robot behavior.
 Language models contain a large amount of information about the real world (Ahn et al., 2022).
-In theory, this should allow the model to exhibit realistic reasoning about robotic behavior.
-However, the models do not have knowledge about particular embodied aspects of a particular robot.
-If we could compare a Scratchpad-like list of intermediate reasoning steps with a list of possible movements of the robot in its environment, then we could prevent the model from suggesting impossible joint movements and actions, and prevent accidents.
-Such an approach has been tried in the Say-can paper (Ahn et al., 2022).
-Say-can learns a value function (Kaelbling et al., 1996) of the behavior of a robot in an environment using temporal difference reinforcement learning Sutton (1988).
-This value function is then combined with prompt-based reasoning by the language model, to constrain it from suggesting impossible or harmful actions.
+In theory, this should allow them to reason realistically about robotic behavior.
+However, the models do not have knowledge about specific embodied aspects of a particular robot.
+If we could compare a Scratchpad-like list of intermediate reasoning steps with a list of possible movements of the robot in its environment, then we could prevent the model from suggesting impossible joint movements and actions, and prevent failures.
+Say-can (Ahn et al., 2022) learns a value function (Kaelbling et al., 1996) of the behavior of a robot and its environment using temporal difference reinforcement learning (Sutton, 1988).
+This value function is combined with prompt-based reasoning by the language model, to constrain it from suggesting impossible actions.
 The goal of Say-can is to ground language in robotic affordances.
-In contrast to Scratchpad, which used supervised learning, the affordance model is learned interactively by reinforcement learning, and then applied using prompt-based learning by the LLM.
-The robot can act as the language model’s hands and eyes, while the language model has high-level semantic knowledge about the task.
-The LLM (Say) provides a task-grounding to find the actions to achieve the high-level goal.
-The learned affordance function (Can) provides a world-grounding to allow what is possible.
-Say-can is evaluated on 101 real-world robotic tasks, such as how to solve tasks in a kitchen environment (see Figure 12).
-Where Say-can learns affordance as a separate function, another approach, Inner-monologue (Huang et al., 2022b) formulates robotic planning directly as part of the language prompt.
-This approach incorporates environmental information into the prompt, linguistically, as an inner monologue.
-As in Say-can, the information comes as feedback from different sources.
-Unlike Say-can, the information of physics and the world is inserted directly into the prompt.
-Inner-monologue consists of many elements: it uses InstructGPT (Brown et al., 2020) for multi-step planning, scripted modules for object recognition, success detection, task-progress scene description, and language-conditioned pick-and-place primitives, similar to CLIPort (Shridhar et al., 2022).
-These elements generate textual descriptions that are used in prompt-based learning.
-Figure 13 gives an example of the working of Inner-monologue.
-The language feedback that is thus generated significantly improves performance on three domains, such as simulated and real table top rearrangement tasks and manipulation tasks in a kitchen environment.
-There are many studies into robotic behavior.
-A recent approach related to Inner-monologue is Chain-of-tools, which proposes a plan-execute-observe pipeline to ground reasoning about tool behavior (Shi et al., 2024a, b).
+In contrast to Scratchpad, which used supervised learning, the affordance model is learned interactively by reinforcement learning, and then applied using prompt-based learning on the LLM.
+The language model has high-level semantic knowledge about the task (Say).
+The learned affordance function (Can) provides an environment-grounding on what is possible.
+Say-can achieves a 31% success rate on 101 real-world robotic kitchen tasks.
+Where Say-can learns affordances as a separate function, Inner-monologue (Huang et al., 2022) formulates robotic planning directly as part of the language prompt, internally.
+The input consists of many elements: textual descriptions from InstructGPT (Brown et al., 2020) for multi-step planning, scripted modules for object recognition, success detection, task-progress scene description, and language-conditioned pick-and-place primitives, similar to CLIPort (Shridhar et al., 2022).
+The language feedback that is thus generated significantly improves performance on three benchmarks, achieving 90% accuracy on simulated and real table top rearrangement tasks and 60% on the kitchen environment.
+There are many other studies into robotic behavior.
+An approach related to Inner-monologue is Chain-of-tools, which proposes a plan-execute-observe pipeline to ground reasoning about tool behavior (Shi et al., 2024a, b).
 This concludes our discussion of the second stage of the reasoning pipeline, evaluation of the reasoning steps.
-5.3 Control of Steps
-The third stage in the reasoning pipeline in Section 2.3 is reasoning control. This stage controls how many sub-steps are generated, and how deep into the future the reasoning chain is generated.
-There are three main approaches: (1) greedy selection, which generates a step and then follows it, (2) ensemble strategy, which generates a set of possible next steps, and (3) a full tree-shaped search which generates multiple options for the step, and follows them multiple steps into the future, traversing a search tree with backtracking, controlling an exponential search space.
-We include reinforcement learning approaches, that interactively learn an optimal policy for such a reasoning space.
-5.3.1 Greedy Selection
-Most earlier works on prompt-based reasoning follow the greedy approach: generate a single prompt with a sequence of steps and follow them.
+3.3. Control of Steps
+The third stage is control.
+This stage controls how many sub-steps are generated, and how deep into the future the reasoning chain is generated.
+There are three main approaches: (3.3.1) greedy selection, which generates a step and then follows it, (3.3.2) ensemble strategy, which generates a set of possible next steps, and (3.3.3) a (reinforcement learning) search which generates multiple options for the steps, traversing a search tree with backtracking, controlling an exponential search space (Xu et al., 2025).
+3.3.1. Greedy Selection
+Most earlier works on prompt-based reasoning follow the greedy approach: generate a single prompt with a single sequence of steps and follow them.
 Among the greedy reasoners are Chain-of-thought, Auto-CoT, and Zero-shot CoT.
 Inner Monologue and Say-Can also use greedy reasoning.
-In Least-to-most prompting (Zhou et al., 2022), the key idea is to break down a complex problem into simpler subproblems and then solve these in sequence, explicitly encoding them in the prompt.
-It is related to Complexity-based prompting.
-In Least-to-most, finding the answer to each subproblem is facilitated by the answers to previously solved subproblems, as in a curriculum (Bengio et al., 2009).
-The authors find that on symbolic manipulation, compositional generalization, and math reasoning, the Least-to-most prompting is capable of generalizing to more difficult problems than those that are given in the prompts.
-Figure 14 illustrates the idea.
-5.3.2 Ensemble Strategy
+In Least-to-most prompting (Zhou et al., 2023), the key idea is to break down a complex problem into simpler subproblems and then solve these in sequence, explicitly encoding them in the prompt, related to Complexity-based prompting.
+In Least-to-most the answers to previously solved subproblems help in finding the answer, as a curriculum (Bengio et al., 2009).
+On symbolic manipulation, compositional generalization, and math reasoning, Least-to-most prompting generalizes well, achieving 99% accuracy on a compositional generalization benchmark.
+3.3.2. Ensemble Strategy
 The second kind of reasoning control is based on an ensemble of (sequences of) reasoning steps.
-The ensemble approach is a well-known technique in machine learning to make a strong learner out of multiple weaker learners (Sagi and Rokach, 2018; Breiman, 2001).
-For most problems, multiple different options for the next step exist.
-When all or some of these are generated and evaluated, then the best result or the consensus result can be reported as the outcome of an ensemble of steps.
-Various approaches have been proposed.
-We already mentioned Self-consistency (Wang et al., 2022b) and Self-verification (Weng et al., 2022) in Section 5.2.1.
-They are popular ensemble approaches to evaluate the results of reasoning steps in prompt learning.
-The greedy single-path decoding used in Chain-of-thought prompting is replaced by sampling a diverse set of reasoning paths, evaluating them, and selecting the most consistent answer.
-In another domain Chain-of-experts builds on Chain-of-thought with a mixture of experts ensemble for complex combinatorial operations research problems (Xiao et al., 2023b).
+For most problems, multiple different options exist for the next step.
+When all or some of these are generated and evaluated, then the consensus result can be reported as the outcome of an ensemble of steps.
+Self-consistency (Wang et al., 2023b) and Self-verification (Weng et al., 2023) (in Section 3.2.1) are popular ensemble approaches to evaluate the results of reasoning steps, in which greedy single-path decoding used in Chain-of-thought prompting is replaced by a diverse set of paths.
+Taking this further, Chain-of-experts uses a mixture-of-experts ensemble for complex combinatorial problems (Xiao et al., 2023b).
 PAL and MathPrompter also use the ensemble approach.
-They generate multiple steps, which are evaluated and whose answer is combined, or the best step is chosen.
-The ensemble approach is a popular approach in LLM-reasoning.
-5.3.3 Reinforcement Learning
-In the greedy approach, a single reasoning path is generated and traversed.
-In reasoning, often multiple valid reasoning steps are possible, but pursuing all possibilities over multiple reasoning steps may lead to an infeasible number of possibilities.
-The third kind of reasoning control is to use a full-fledged controller that can traverse a tree, or even perform reinforcement learning to do so (Sutton and Barto, 2018; Kaelbling et al., 1996; Plaat, 2022).
-This group of control approaches enables the most elaborate control of the reasoning process, and is used by many works, as we will see.
+The ensemble approach is popular in reasoning with LLM.
+3.3.3. Reinforcement Learning
+In reasoning, often multiple valid steps are possible, but pursuing all possibilities over multiple trajectories may lead to an infeasible number of possibilities.
+The third kind of reasoning control is to use a full-fledged controller that can traverse a tree, or perform reinforcement learning to do so (Sutton and Barto, 2018; Kaelbling et al., 1996; Plaat, 2022).
 When decomposing the problem, multiple alternative steps are generated that can be searched multiple steps into the future.
 Then, backtracking can be performed, allowing alternative steps to be tried.
-Where greedy and ensemble processes can be controlled with a prompt by the LLM, this third group is more complex, and an external algorithm is used to control the reasoning process.
-The external algorithms call the LLM as a subroutine prompting it to perform its tasks.
-This allows more complex reasoning control, but we are no longer performing prompt-based self-reasoning; control has been given to an algorithm that is external to the LLM and external to prompt-learning.
+Where greedy and ensemble processes can be controlled with a prompt by the LLM, this third category is more complex, and an external algorithm is used to control the reasoning process.
+The external algorithms call the LLM as a subroutine prompting it to perform requested tasks.
+The external algorithm allows more complex reasoning control, but we are now beyond prompt-based self-reasoning: control has been given to an algorithm that is external to the LLM and external to prompt-learning.
 We start our discussion of control strategies with depth-first and breadth-first search, then go to beam search, and then to full reinforcement learning.
-Breadth first search
 A complex reasoning space can be traversed with a search algorithm.
-Tree-of-thoughts includes a search algorithm to dynamically follow different reasoning steps (Yao et al., 2024).
-When one reasoning path has been traversed, a search algorithm can backtrack, and try an alternative path.
-The paper describes both a breadth-first-search and a depth-first-search controller.
-The evaluation part in Tree-of-thoughts is performed with a prompt by the LLM.
+Tree-of-thoughts (Yao et al., 2024) uses breadth-first or depth-first search to dynamically follow different reasoning steps.
+The evaluation part in Tree-of-thoughts is performed with a prompt to the LLM.
 Together, the trio of generation, evaluation, and control allow systematic exploration of the space of reasoning steps with look-ahead and backtracking.
-The authors compare their approach to Chain-of-thought and Self-consistency.
-Chain-of-thought builds a reasoning out of a path of thoughts, Self-consistency creates an ensemble of thoughts, and Tree-of-thoughts constructs a tree structure.
-Figure 15 illustrates the different reasoning structures.444A similarly named approach is Graph-of-thoughts (Besta et al., 2024).
+The authors compare their approach to Chain-of-thought and Self-consistency on the Game of 24, Creative writing, and Mini crossword, achieving an accuracy of 74% on a Game of 24 task.
+The other tasks are evaluated qualitatively.
+Figure 5 illustrates the different reasoning structures.333A similarly named approach is Graph-of-thoughts (Besta et al., 2024).
 Graph-of-thoughts allows more general reasoning graphs, providing a formal framework, where the different elements can then be specified manually.
-Another approach, Buffer-of-thoughts (Yang et al., 2024), goes a step further towards meta-reasoning.
+Another approach, Buffer-of-thoughts (Yang et al., 2024), goes a step towards metareasoning (Gao et al., 2024).
 It introduces a meta-buffer that stores high-level thought-templates.
-These universal thought-templates are derived from a variety of tasks.
+These thought-templates are derived from a variety of tasks.
 Figure 5 compares the Buffer-of-thoughts approach to other approaches such as Chain-of-thought and Tree-of-thoughts.
-Buffer-of-thoughts outperforms other methods in puzzles such as Game of 24 and checkmating.
+Buffer-of-thoughts outperforms other methods in puzzles such as Game of 24 (by 11%) and checkmating (by 51%).
 Thought templates are related to metacognition (thinking about thinking), which is further discussed in Section 4.3.3.
-Beam search
-A related search method is Beam-search.
-Beam-search-for-reasoning (Xie et al., 2024) focuses on control of the space of possible reasoning paths.
-In some reasoning problems, this space can be very large.
-Beam-search solves this challenge by searching only a promising part of this space.
+A related search method is Beam-search-for-reasoning (Xie et al., 2024).
+When the space of possible reasoning paths is large, Beam-search searches a promising part of this space.
 It uses self-evaluation to control exploration and to evaluate (decode) reasoning steps.
-Figure 17 shows how Beam-search self-evaluation is used in multi-step reasoning.
 Beam search uses Program-aided-language models for math word problems (Gao et al., 2023).
 Using a Codex backbone (Chen et al., 2021), it surpasses the few-shot baselines by 6.34%, 9.56%, and 5.46% on the GSM8K, AQuA, and StrategyQA benchmarks, respectively.
-Reinforcement learning
-Reinforcement learning (RL) methods are another step in the sophistication of optimization algorithms.
-RL learns by interactive sampling, improving its policy based on rewards from the environment (Sutton and Barto, 2018).
-To use reinforcement learning, the reasoning problem must be formulated as a Markov Decision Process: the agent-algorithm creates a prompt (an action), to sample a step () and get an answer (state, reward) from the environment-model (see Figure 18).
-The answer can then be used to improve the prompt (next action), just like reinforcement learning uses rewards to improve its policy of best actions for each state.
-The approaches that use reinforcement learning do so in the form of an external algorithm.
-No prompt has been created that performs RL by itself.
-Progressive-hint-prompting (PHP) uses reinforcement learning to interactively improve prompts (Zheng et al., 2023). Figure 19 illustrates the approach.
-PHP is an external algorithm that calls the LLM with dynamic prompts, using previously generated answers as hints to progressively prompt the LLM toward the correct answers.
-It works as follows: (1) given a question (prompt), the LLM provides a base answer, and (2) by combining the question and answer, the LLM is queried and we obtain a subsequent answer.
-We (3) repeat operation (2) until the answer becomes stable, like a regular policy-optimizing reinforcement learning algorithm.
+Reinforcement learning is another step in the sophistication of optimization algorithms.
+It learns by interactive sampling, improving its policy based on rewards from the environment (Sutton and Barto, 2018).
+To use reinforcement learning, the reasoning problem is formulated as a Markov Decision Process: the agent-algorithm creates a prompt (an action), to sample a step () and get an answer (state, reward) from the environment-model.
+The answer can then be used to improve the prompt (next action), using the rewards to improve its policy of best actions for each state.
+The approaches that use reinforcement learning also do so in the form of an external algorithm.
+Progressive-hint-prompting (PHP) uses reinforcement learning to interactively improve prompts (Zheng et al., 2023).
+PHP calls the LLM with dynamic prompts, using previously generated answers as hints, to progressively prompt the LLM towards the correct answers.
+It works as follows: (1) given a question (prompt), the LLM provides a base answer, and (2) by combining the question and answer, the LLM is queried and obtains a subsequent answer.
+We (3) repeat operation (2) until the answer becomes stable, as a regular policy-optimizing reinforcement learning algorithm.
 The authors have combined PHP with Chain-of-thought and with Self-consistency.
-Using GPT-4, state-of-the-art performance was achieved in grade school math questions (95%), simple math word problems (91%) and algebraic question answering (79%).
+Using GPT-4, state-of-the-art performance was achieved on grade school math questions (95%), simple math word problems (91%) and algebraic question answering (79%).
 Another approach that is motivated by improving answers from feedback, is Self-refine (Madaan et al., 2023).
-In this method, initial outputs from LLMs are improved through iterative feedback and refinement.
-Like PHP, the LLM generates an initial output and provides feedback for its answer, using it to refine itself, iteratively.
-Figures 20 and 21 illustrate the approach.
-Self-refine prompts the LLM in three ways: (1) for initial generation, (2) for feedback, and (3) for refinement.
-Note that Self-refine follows a greedy reasoning chain, learning from feedback.
-Self-refine has been used with GPT-3.5 and GPT-4 as base LLMs, and has been benchmarked on dialogue response generation (Askari et al., 2024), code optimization, code readability improvement, math reasoning, sentiment reversal, acronym generation, and constrained generation, showing substantial improvements over the base models.
-Another approach that combines reinforcement learning and LLMs is ReAct (Yao et al., 2022).
-Most works so far have focused on reasoning by the LLM, not on actions by an agent.
-A key element of reinforcement learning is that it learns a policy for an environment.
+Like PHP, the LLM generates an initial output and provides feedback for its answer, using the LLM to refine itself, iteratively.
+Self-refine prompts the LLM in three ways: (1) for initial generation, (2) for feedback, and (3) for refinement, following a greedy reasoning chain.
+Self-refine has been used with GPT-3.5 and GPT-4 as base LLMs, and has been benchmarked beyond math word problems on dialogue response generation (Askari et al., 2024), code optimization, code readability improvement, math reasoning, sentiment reversal, acronym generation, and constrained generation, showing substantial improvements over the base models (typically around 30 percentage points, see Table 2).
+Another approach that combines reinforcement learning and LLMs is ReAct (Yao et al., 2023).
+Most works focus on reasoning by the LLM, not on actions by an agent.
 The goal of ReAct is to combine progress in reasoning with action plan generation.
-(Or, to put it differently, most approaches use RL to improve LLM-reasoning, ReAct uses LLMs to improve RL agent policies.) ReAct uses Chain-of-thought prompt-learning as part of an RL framework that also uses external knowledge sources (Wikipedia) and finetuning, for error reduction, grounding, and for reducing hallucination.
+(Or, to put it differently, other approaches use RL to improve LLM-reasoning, ReAct uses LLMs to improve RL agent policies.) ReAct uses Chain-of-thought prompt-learning as part of an RL framework that also uses external knowledge sources (Wikipedia) and finetuning; for error reduction, grounding, and for reducing hallucination.
 The framework allows hand-written prompts.
-Figure 22 shows four different prompting strategies.
-On two interactive decision making benchmarks (ALFWorld and WebShop), ReAct outperforms imitation and reinforcement learning methods by an absolute success rate of 34% and 10% respectively, while being prompted with one or two in-context examples.
+On two interactive decision making benchmarks (Alfworld and WebShop), ReAct outperforms imitation and reinforcement learning methods by an absolute success rate of 34% and 10% respectively.
 The ReAct work has been developed further.
-Reflexion (Shinn et al., 2024) is built on top of ReAct.
-The goal is to create AI agents that learn by reflecting on failures and enhancing their results, much like humans do.
+Reflexion (Shinn et al., 2024) creates AI agents that learn by reflecting on failures and enhance their results, much like humans do.
 Reflexion uses three language models: actor, evaluator, and reflector.
 It works as follows: (1) an actor generates text and actions, (2) an evaluator model scores the outputs produced by the actor, and (3) a self-reflection model generates verbal reinforcement cues to assist the actor to self-improve (see Figure 6).
-For the actor, Chain-of-thought (Wei et al., 2022b) and ReAct (Yao et al., 2022) can be used.
+For the actor, Chain-of-thought and ReAct can be used.
 Reflexion is evaluated on decision-making, reasoning, and coding tasks.
 Improvements of 10-20 percentage points are reported.
-Figure 24 shows three different prompting applications.
+The reinforcement learning approaches that we discussed so far—React, Self-refine, Tree-of-thoughts, Buffer-of-thoughts, Reflexion—use external algorithms to manage state for the prompt improvement loop.
+A more elegant solution would be to perform reinforcement learning fully in-context, within-prompt.
+Indeed, Krishnamurthy et al.
+(2024) explicitly asked the question whether LLMs can explore in context.
+This is the goal of the Algorithm-of-thoughts (AoT) approach (Sel et al., 2023).
+Following work by Lee et al.
+(2023); Wang et al.
+(2024c); Demircan et al.
+(2024) that showed that transformer architectures can be pretrained and finetuned to perform in-context reinforcement learning, AoT aims to do so in general LLMs such as GPT-4, Claude and Gemini 1.5.
+They achieve results comparable to Tree-of thoughts on GSM8K, StrategyQA, and Crosswords.
+Achieving these results with in-context RL is a promising result for in-context learning.
+Other search-like in-context algorithms are studied by Schultz et al.
+(2024) and Kempinski et al.
+(2025) (Game-of-thoughts).
+Further work in games finds that LLMs struggle with the knowing-doing gap and the difference between using and mentioning game concepts (Paglieri et al., 2024; van Wetten et al., 2025; Cloos et al., 2024).
 To conclude this overview of reinforcement learning approaches, we discuss an application in the games domain.
-Voyager (Wang et al., 2023) is an agent for the game of Minecraft that uses an iterative prompting mechanism that generates code for embodied control.
-The mechanism includes Self-verification (Shinn et al., 2024).
-The agent has a skill library and an automatic curriculum to maximize exploration.
-Voyager interacts with GPT-4 through prompts.
-The goal of Voyager’s prompts is to discover as many diverse items in Minecraft as possible, a form of novelty search (Eysenbach et al., 2018).
-Voyager performs well, it shows in-context lifelong learning capability and reaches high scores by acquiring many tools (see Figure 25).
-6 Discussion
-We have reviewed approaches for prompt-based reasoning by LLMs, highlighting techniques that have achieved a breakthrough in reasoning performance.
-It is time for reflection on limitations in the approaches, suggesting promising areas of future work.
-First we discuss issues concerning hallucination, faithful reasoning, and scaling.
+Voyager (Wang et al., 2024a) is an agent for the game of Minecraft that uses an iterative prompting mechanism that generates code for embodied control.
+The agent includes self-verifcation an a skill library to maximize exploration.
+The goal of is to discover diverse items in Minecraft, a form of novelty search (Eysenbach et al., 2019).
+Voyager performs well, reaching high scores by acquiring many tools (see Figure 7) 15 times faster than the baseline.
+The applications of reinforcement learning in LLM reasoning are many, and the connections run deep (Pternea et al., 2024).
+Wang et al.
+(2024b) use the similarity between RL timesteps and LLM reasoning steps to jointly train a value function together with the LLM policy by optimizing the soft Bellman equation, achieving 85% accuracy on GSM8K and 81% on Alfworld.
+Guo et al.
+(2025); Du et al.
+(2025) replace supervised finetuning by reinforcement learning, integrating it with reasoning, achieving large efficiency gains in the training pipeline in the DeepSeek r1 and Kimi models.
+4. Discussion
+We have reviewed many reasoning approaches.
+It is now time to reflect on the approaches, to discuss limitations, and to look for promising areas of future work.
+First we discuss how to choose an approach for an application.
+Next, we discuss issues concerning hallucination, faithful reasoning, and scaling.
 Then we discuss what LLMs can and cannot do.
-Then, we highlight connections with sequential decision processes and metacognition, and end with a research agenda.
-6.1 Hallucination, Faithfulness and Scaling
-Most works on reasoning in LLMs are experimental in nature.
-The success of in-context learning and Chain-of-thought reasoning is attracting the attention of work providing deeper insight into the reasoning processes in language models.
+Then, we highlight connections with sequential decision processes, metacognition, and programming languages, and we end with a research agenda.
+4.1. Matching Methods and Applications
+This survey has discussed different approaches, applications, and benchmarks.
+It is often difficult to directly compare benchmark results between individual papers, due to differences in measurement setup.
+For example, Kamoi et al.
+(2024) critically compare self-reflection studies.
+We can provide, however, general guidance on what approach to use for different types of applications.
+For simple reasoning situations, that require a single sequence of reasoning steps, Chain-of-thought is typically used.
+When the error rate is high, verification methods such as Self-consistency are popular.
+Indeed, this combination is used in many modern reasoning LLMs.
+For combinatorial puzzles and games that would traditionally require backtracking, tree-search or reinforcement learning, solution methods need to be able to keep track of the enumeration state.
+However, an LLM without an external algorithm often fails to keep track of this state correctly between calls to the model.
+An approach such as Tree-of-thoughts, or ReAct, should be considered.
+Here, the state is managed externally, and the LLM is called for application-dependent functional processing.
+Various in-context reinforcement learning methods are being developed where state is managed by the LLM, as in, for example, Algorithm-of-thoughts.
+When performance is not sufficient, finetuning or pretraining for the domain at hand can be used.
+For complex problems, where a sequence of steps or external control fail to perform well, other reasoning approaches are needed.
+We note that LLMs are good at generating computer code for many known problems (Chen et al., 2021).
+Hence, when the preceding approaches do not work, the LLM can be asked to generate a program for the problem.
+In this approach, the LLM acts as a coding agent for its user.
+The LLM generates a program or a sequence of commands, which are then executed by an external tool, such as an interpreter, a robot, a planner, or a logic solver.
+In robotics, vision-language-action models have achieved impressive results (Kim et al., 2024), but also strong results in logistics (Bohnet et al., 2024), finance and medicine are reported (Plaat et al., 2025).
+4.2. Hallucination, Faithfulness and Scaling
+Reasoning by LLMs is not error-free, and many studies aim to provide deeper insight into the reasoning processes in language models.
 Saparov and He (2023) introduce a synthetic question/answer dataset designed to evaluate the reasoning abilities of LLMs.
@@
 In another study, Wang et al.
-(2022a) also aim to increase our understanding of how Chain-of-thought works.
-The authors find that it continues to work even with invalid steps in the reasoning chain.
-They also find that the order of the reasoning steps is important for good results.
+(2023a) aim to increase our understanding of how Chain-of-thought works.
+The authors find that the order of the reasoning steps is important.
 Prompts should be relevant to the question, and coherent (steps should be in the correct order).
 Jin et al.
-(2024) study the impact of reasoning step length on LLMs, finding a strong positive correlation between the length of the prompt and reasoning abilities.
-These works highlight ways in which LLM-reasoning can see things that are not there.
-Next, we discuss works on failure modes of the Chain-of-thought approach, studying whether the reasoning of the LLM is faithful, or that it gives the right answer for the wrong reason.
-6.1.1 Faithfulness
-Chain-of-thought and other approaches prompt a language model to take certain steps to solve the problem that the prompt specifies.
-One can ask the question, whether those steps are indeed the steps that the model has followed (faithful reasoning) or whether it took another road to arrive at the correct answer (unfaithful reasoning).
-A few studies measure the faithfulness of reasoning by LLMs.
+(2024) also study the impact of reasoning step length on LLMs, again finding a strong positive correlation between the length of the prompt and reasoning abilities.
+Next, we discuss works on errors in the Chain-of-thought approach, studying whether the reasoning of the LLM is faithful, or that it gives the right answer for the wrong reason.
+4.2.1. Faithfulness
+Chain-of-thought approaches prompt a language model to take certain steps to solve the problem that the prompt specifies.
+One can ask the question, whether those steps are indeed the steps that the model has followed (faithful reasoning) or whether it took another road to arrive at the same answer (unfaithful reasoning).
+A few studies measure the faithfulness of reasoning with LLMs.
 Lanham et al.
@@
 They find a wide variation of post-hoc reasoning, with a tendency of larger models to be less faithful.
-Like regular LLMs, when not properly grounded, (Chain-of-thought) reasoning suffers from hallucination.
+Like regular LLMs, when not properly grounded, (Chain-of-thought) reasoning suffers from hallucination (Huang et al., 2025).
 Another study adds deliberate bias to the prompt.
@@
 They find that a bias towards wrong answers can cause significant drops in accuracy, and that models frequently generate Chain-of-though explanations rationalizing wrong answers.
-The authors further note that, insofar as language models are trained on human-written explanations, that explanations may be incomplete or wrong.
-Human explanations may omit crucial steps of the causal chain, may provide an unfaithful account of the human reasoning process, or may be aimed at convincing others, instead of providing the true causes of a decision.
 To address issues of faithfulness, Lyu et al.
-(2023) propose Faithful-chain-of-thought.
+(2023); Xu et al.
+(2024a) propose Faithful-chain-of-thought.
 This approach involves two stages.
@@
 For the symbolic language, Python, Datalog, or PDDL is suggested.
-Faithfulness studies tell us more about how models reason.
-Further surveys on this topic are Mondorf and Plank (2024); Chuang et al.
-(2024); Luo et al.
-(2023); Paul et al.
-(2024),
-6.1.2 Scaling
+Another approach, mechanistic interpretability, studies methods to target individual representations inside the LLM, to see if the expected behavior occurs in practice (Rai et al., 2024; Bereska and Gavves, 2024; Chen, 2025).
+4.2.2. Scaling
 The emergent abilities of LLMs have prompted research into the nature of scaling and reasoning with LLMs, and, specifically, how reasoning capabilities can be transferred to smaller language models.
-Scaling laws of LLMs are an active area of study, see for example Kaplan et al.
-(2020); Henighan et al.
-(2020); Hoffmann et al.
-(2022).
-Given the computational cost of LLMs, there is much interest in transferring knowledge to small language models.
+Scaling laws of LLMs are an active area of study, see for example (Kaplan et al., 2020; Henighan et al., 2020; Hoffmann et al., 2022).
+Distillation of reasoning to smaller models can work surprisingly well in situations with more explicit instructions, and given the computational cost of training LLMs, there is much interest in transferring knowledge to small language models.
 Comprehensive surveys on knowledge distillation are Xu et al.
-(2024); Gu et al.
+(2024b); Gu et al.
 (2023).
@@
 (2023) have studied reasoning in small language models, using a student model that learns from a teacher model, by finetuning.
-Another study related to Self-taught-reasoner (Li et al., 2022a) focuses on explanation in small language models, achieving similar results.
-Other works focus on prompt distillation for retrieval Dai et al.
-(2022), recommendation (Li et al., 2023), distillation to embodied agents of Chain-of-thought reasoning (Choi et al., ), and distillation of LLM graph reasoning (Zhang et al., 2024).
-Distillation of reasoning to smaller models can work surprisingly well in situations with more explicit instructions.
-Distillation is also proposed for bringing results of System 2 reasoning to System 1 Yu et al.
-(2024), which brings us to the topic of metacognition (see Section 6.2.3).
-6.2 Limitations: What LLMs Can and Cannot do
+Other works focus on prompt distillation for retrieval (Dai et al., 2023), recommendation (Li et al., 2023a), embodied agents (Choi et al., ), and LLM graph reasoning (Zhang et al., 2024).
+Another study related to Self-taught-reasoner focuses on explanation in small language models, also achieving good results of knowledge transfer (Li et al., 2024).
+4.3. Limitations: What LLMs Can and Cannot do
 The capabilities of LLMs are impressive.
-LLMs can be seen as large text-based surrogate models of the world (or the world how we describe it on the internet), and thus allow us to reason in a way that we can understand about a large variety of contexts and problems.
-Reasoning tasks, such as math word problems, were one of the capabilities that LLMs could not achieve, until recently.
-Let us look more closely at what language models can and cannot do.
-6.2.1 What Can LLMs Do?
-With the right prompt LLMs are able to solve many of the problems in reasoning grade school math word benchmarks.
-Prompt-based learning is able to perform reasoning tasks such as math word problems, robotic movement, and Python code generation, at inference time, without expensive parameter training.
-We note that a simple taxonomy of generate-evaluate-control is able to describe the structure of the current LLM reasoning literature well.
-Furthermore, the accuracy of the reasoning chains can be improved with ensemble methods, or self-verification.
+LLMs can be seen as large text-based surrogate models of the world (or the world how we describe it on the internet), and thus allow reasoning about a large variety of contexts and problems.
+Reasoning, such as in math word problems, were one of the capabilities that LLMs could not achieve, until recently.
+Let us look more closely at what language models currently can and cannot do.
+4.3.1. What Can LLMs Do?
+With the right prompt, LLMs are able to solve many of the problems in grade school math word benchmarks and beyond.
+Prompt-based learning is able to perform reasoning tasks such as math word problems, robotic movement gaming, and Python code generation, at inference time, without expensive parameter training.
+A taxonomy of generate-evaluate-control is able to describe the structure of the current LLM reasoning literature.
+Furthermore, the accuracy of the reasoning chains can be improved with ensemble methods, and self-verification.
 Hallucination can be reduced by grounding the model with external models, such as for robotic affordances, and information retrieval from search engines and Wikipedia.
-Going a step further, using external control algorithms (such as search or RL) as scaffolding, dynamic prompts can use the LLMs to perform complex and interactive reasoning patterns.
+Going a step further, using external control algorithms as scaffolding (such as search or reinforcement learning), dynamic prompts can use the LLMs to perform complex and dynamic reasoning patterns.
 Note that the reasoning control is now two layers away from the core LLM: an external control algorithm, on top of in-context-learning, dynamically generating prompts for the LLM.
 This is reasoning with prompts with LLMs, not by.
-At this point, it is interesting to note the confluence of the two schools of classical artificial intelligence (AI), symbolic and connectionist.555Reasoning and planning have been studied since the start of artificial intelligence, starting with logic and reasoning (Newell and Simon, 1961), search algorithms in puzzles and board games (Korf, 1999; Plaat, 2020), robot planning (Fikes and Nilsson, 1971), classical machine learning such as decision trees and support vector machines (Flach, 2012; Breiman, 2001; Cortes and Vapnik, 1995), through knowledge representation and the semantic web (Van Harmelen et al., 2008).
-Ever since the success of the connectionist approach LeCun et al.
-(2015); Goodfellow et al.
-(2016) (deep learning, including LLMs) researchers have tried to join the two approaches.
-Search and reinforcement learning are rooted in the symbolic AI tradition, while LLMs are rooted in the connectionist tradition.
+At this point, it is interesting to note the confluence of the two schools of classical artificial intelligence, symbolic and connectionist.444Reasoning and planning have been studied since the start of artificial intelligence, starting with logic and reasoning (Newell and Simon, 1961), search algorithms in puzzles and board games (Korf, 1999; Plaat, 2020), robot planning (Fikes and Nilsson, 1971), classical machine learning such as decision trees and support vector machines (Flach, 2012; Breiman, 2001; Cortes and Vapnik, 1995), through knowledge representation and the semantic web (Van Harmelen et al., 2008).
+Ever since the success of the connectionist approach (LeCun et al., 2015; Goodfellow et al., 2016) (deep learning, including LLMs) researchers have tried to join the two approaches.
+Search and reinforcement learning are rooted in the symbolic tradition, while LLMs are rooted in the connectionist tradition.
 The literature in this survey combines the two traditions.
 High performance reasoning is created with a (symbolic) searcher/learner on top of a (connectionist) LLM.
-In other fields similar combinations can be seen (for example, AlphaFold Bryant et al.
-(2022); Jumper et al.
-(2021) and retrosynthesis of molecules Segler et al.
-(2018)).
-The LLM helps ground symbolic reasoning methods in language; symbolic methods help create prompts that let the LLM perform reasoning.
-How the two traditions will continue to improve eachother, we will see in further research.
-We note that benchmarks such as GSM8K have been central for the progress in the field, and that while reasoning started with math word problems, the field has extended to robotics, autonomous agents, games, and most emphatically computer code.
+In other fields similar combinations can be seen (for example, AlphaFold (Bryant et al., 2022; Jumper et al., 2021) and retrosynthesis of molecules (Segler et al., 2018)).
+The LLM helps ground symbolic reasoning methods in language; symbolic methods help create prompts that let the LLM perform dynamic reasoning.
+We note that benchmarks such as GSM8K have been central for the progress of the field, and that while reasoning started with math word problems, the field has extended to robotics, autonomous agents, games, and most emphatically computer code.
 Formal languages play an important role in the intermediate multi-step reasoning chains.
-A side effect from the work on reasoning is the emergence of a new few-shot learning approach for sequential decision-making processes (SDP)(Littman, 1996).
-Traditionally these processes are solved with reinforcement learning (such as DQN Mnih et al.
-(2015), PPO (Schulman et al., 2017) and SAC Haarnoja et al.
-(2018)), achieving good results, but suffering from high sample complexity for larger problems Plaat et al.
-(2023).
-The emergence of few-shot in-context learning for solving SDPs opens a research avenue to find out what SDPs few-shot prompt-learning will be able to solve.
-6.2.2 What Can LLMs Not Do?
+4.3.2. What Can LLMs Not Do?
 Now that grade school math word problems are largely solvable, harder reasoning benchmarks in other domains are appearing (Ahn et al., 2024).
-Another line of research argues that LLMs cannot reason, providing examples where LLMs fail, and discussing potential reasons.
-Berglund et al.
-(2023) show that LLMs can fail to generalize in surprising ways.
-They provide the example that if a model is trained to report that ”Valentina Tereshkova was the first woman to travel to space”, it will not automatically be able to answer the question, ”Who was the first woman to travel to space?” pointing to a lack in semantic understanding of LLMs.
-Other work suggests that results are less generalizable and transferable than often assumed, showing how base-10 arithmetic skills do not transfer to base-9 arithmetic problems Wu et al.
-(2024).
-The question which problems LLMs can and cannot solve will continue to motivate researchers.
+Most of the reasoning capabilities exhibited by LLMs are due to the representational powers of the transformer architecture, and how in-context learning is able to harness them.
+Prompt engineering and prompt control play a crucial role in the kind of reasoning that we have seen in the papers.
+Models can be instructed to write their own reasoning prompts; however, such Auto-GPT or Auto-CoT prompts need careful evaluation, verification, and grounding in the real world, to prevent degeneration into a hallucinatory world of their own.
+Models can also be instructed to interact with the world, and become the tool of external scaffolding that evaluates, controls and improves the prompts (Plaat et al., 2025).
+Some of what we experience as reasoning by the LLM, is controlled by the prompt or the scaffolding algorithm.
+Studies into in-context reinforcement learning aim to answer the question if prompt learning is able get the LLM to create a prompt to exhibit dynamic reasoning by itself (Lee et al., 2023; Schultz et al., 2024; Demircan et al., 2024; Paglieri et al., 2024).
+Some studies on symbolic planning are critical on the abilities of LLMs (Valmeekam et al., 2023), and show examples of planning failures, arguing that LLMs are better used to improve heuristic elements of traditional planners, such as PDDL (Kambhampati et al., 2024), to strengthen traditional symbolic planning approaches.
 Other works study the dangers of the size of LLMs.
@@
 Finally, to prevent putting too much trust in the outcome of LLMs, we should understand their failure modes better, such as the well-publicized problems of hallucination (inventing facts that look right but are not).
-Most of the reasoning capabilities exhibited by LLMs are due to the great representational powers of the transformer architecture, and how in-context learning is able to harness them.
-Prompt engineering and prompt control play a crucial role in the kind of reasoning that we have seen in the papers.
-Models can be instructed to write their own reasoning prompts; however, such Auto-GPT or Auto-CoT prompts need evaluation, verification, and grounding in the real world, to prevent degeneration into a hallucinatory world of their own.
-Models can also be instructed to interact with the world, and become the tool of external scaffolding that evaluates, controls and improves the prompts.
-Some of what we experience as reasoning by the LLM, is controlled by the prompt or the scaffolding algorithm.
-It is an open question if prompt learning is able get the LLM to create a prompt to exhibit non-trivial reasoning by itself.
-From the symbolic planning field there is also a critical view on the reasoning and planning abilities of LLMs (Valmeekam et al., 2023) giving examples of planning failures.
-They argue that LLMs can be used instead to improve heuristic elements of traditional planners, such as PDDL (Kambhampati et al., 2024), to strengthen traditional symbolic planning approaches.
+Here, mechanistic interpretability can be used to explore LLM representations and understand where they go wrong (Bereska and Gavves, 2024; Rai et al., 2024; Chen, 2025).
 Some of the names of the approaches surveyed in this paper are suggestive of self-awareness and self-reflective capabilities.
-True self-reflection, or metacognition, is still largely outside the capabilities of current LLMs.
+True (human) self-reflection, or metacognition, is still largely outside the capabilities of current LLMs.
 LLMs can be prompted to reason, to take small steps, to self-evaluate, and their search process can be controlled by an external algorithm.
-The self-reflective type of “intelligence” is written into the prompt by the prompt engineer or the interactive algorithm.
+The self-reflective type of “intelligence” is written into the prompt by the prompt engineer or the control algorithm.
 We are unaware of any LLM that has been made to reflect on, or even control, its reasoning processes, controlling how many reasoning steps it should take, or limiting its reasoning once the answer had become good enough.
 True self-reflection remains future work, although some steps have been taken, as we will discuss next.
-6.2.3 Reasoning towards Metacognition
+4.3.3. Reasoning towards Metacognition
 Human thought exhibits the ability to reason about self, we are able to think about our own thinking processes.
-Metacognition studies these topics (Veenman et al., 2006).
-Prompted by the success of Chain-of-thought and the works that we have surveyed, metacognition has also been studied in the context of LLMs (Toy et al., 2024).
+Metacognition studies this phenomenon (Veenman et al., 2006).
+Prompted by the success of Chain-of-thought and related works, metacognition has also been studied in the context of LLMs (Toy et al., 2024).
 Many reasoning approaches highlight self-reflective aspects in their names and in how they work.
@@
 Inspired by Chain-of-thought and Self-consistency, they create manually designed prompts to increase the understanding of language models.
-Figure 26 illustrates the relation between metacognitive human thought processes and metacognitive LLM prompting.
 Another work, again inspired by Chain-of-thought and Self-consistency, connects psychology and LLMs.
@@
 Interestingly, the authors find that the skill exemplar repository that is trained on a strong model (GPT-4), also down-translates to a weak model (GPT-3).
-The performance of the weak model benefits from the skill-name-enhanced prompts.
-The connection between reasoning in LLMs and full-blown metacognitive reasoning is in its early stages. Exciting future research may appear.
-6.3 Research Agenda
-At the end of this discussion, we present promising topics for future work.
+The performance of the weaker model benefits from the skill-name-enhanced prompts.
+Begus et al.
+(2025) propose a program for the study of metalinguistic abilities of LLMs.
+LLMs struggle with analogies (Lewis and Mitchell, 2024; Mitchell and Hofstadter, 1990) and analogical reasoning.
+Metacognitive reasoning with LLMs is still in its early stages.
+4.4. LLMs that Generate Computer code
+A persistent thread in the results in this survey is that LLMs are good at generating formal languages, such as PDDL, logic, math equations, and Python.
+The agentic approach in which LLMs are combined with external tools such as interpreters or debuggers often yields good performance (Plaat et al., 2025).
+LLMs can be used in different modes, which we will now discuss.
+When LLMs are used in direct mode, a prompt is provided with an instruction, and an answer. This is the mode in which most casual users use an LLM.
+LLMs can also be used in algorithmic mode, or Chain-of-thought mode, where an extra element is added to the prompt that gives an example of which steps to take to arrive at the answer (Sel et al., 2023).
+Lee et al.
+(2023); Wang et al.
+(2024c); Schultz et al.
+(2024) show that the transformer architecture can be trained such that search and reinforcement learning algorithms can be executed in-context.
+(LLMs pretrained on natural language are typically not good at in-context reinforcement learning (Ruoss et al., 2024; Paglieri et al., 2024)).
+We have seen that LLMs are not good at preserving state between calls (Kempinski et al., 2025).
+External algorithms can be added to manage the state, that call the LLM in an external mode (Yao et al., 2024, 2023).
+Tree-of-thoughts, and similar approaches, use the LLM in a stateless functional way, with the prompt containing the instructions, not unlike a natural language variant of a functional programming language such as LISP or Haskell.
+Finally, an LLM can be used in code generation mode.
+The prompt specifies a problem, and the LLM is asked to generate computer code (such as Python or PDDL).
+The code may contain variables that manage the state correctly.
+The code can be executed directly, or first be checked and improved by a programmer.
+We expect that the use LLMs as code generators, and prompt-engineering as natural-language-programming, will continue to grow.
+4.5. Research Agenda
+At the end of this discussion, we list promising topics for future work.
 Reasoning with LLMs is an active field of research.
-It brings together elements of symbolic reasoning, connectionism, natural language, autonomous agents, and affective reasoning (Broekens et al., 2023) with the promise of artificial general intelligence.
-For the future, the surveyed works point in the following directions.
-First we discuss topics for the field of LLM-reasoning itself, then we discuss more general machine learning topics that are important for progress in LLM-reasoning, and finally we discuss more longer term, fundamental topics.
-Specific research topics for reasoning with LLMs are:
-Control and prompt-learning—Search control beyond greedy search is implemented as an external algorithm.
-Is it possible to incorporate all stages of the reasoning pipeline into an interactive prompt?
-Can we make a prompt that performs dynamic search-like step control without external scaffolding?
+It brings together elements of symbolic reasoning, connectionism, natural language, autonomous agents, affective reasoning (Broekens et al., 2023) and metacognition.
+First we discuss current topics for the field of LLM-reasoning itself, then we discuss more general machine learning topics that are important for progress in LLM-reasoning, and finally we discuss more longer term, fundamental topics.
+In Context Reinforcement Learning—Search control beyond greedy search is often implemented as an external reinforcement learning algorithm.
+Is it possible to incorporate the control stage of the reasoning pipeline into one static prompt, as in-context reinforcement learning (ICRL)?
+Studies indicate that models can be trained for decision making (Lee et al., 2023; Demircan et al., 2024), but general LLMs have not yet been shown to perform ICRL well (Schultz et al., 2024; Paglieri et al., 2024; Ruoss et al., 2024; van Wetten et al., 2025).
 Code—Progress in reasoning using formal languages and computer code has been quite promising.
@@
 A challenge is the grounding of the reasoning process in the environment.
-How can we help LLMs to actively find new information when the reasoning outcome is uncertain?
-Is retrieval augmented generation the future?
-Is the future of the reasoning-LLM a search engine (Verberne, 2024)?
+Retrieval augmented generative methods (RAG) can help LLMs to actively find new information when the reasoning outcome is uncertain.
+RAG, search, and agentic LLMs, are also active areas of research (Dai et al., 2023; Verberne, 2024; Plaat et al., 2025).
 Generic topics in machine learning that also influence prompt-based reasoning research are:
-Benchmarks—Progress in LLMs is governed by the availability of the right benchmarks.
-The current favorite is GSM8K, for grade school math.
-As the field progresses, other benchmarks will become prevalent: benchmarks with more difficult tasks, and benchmarks for other applications in autonomous agents and robotics.
+Benchmarks—Progress in LLMs depends on the availability of the right benchmarks.
+As the field has progressed beyond math word problems, other benchmarks become prevalent, with more difficult and diverse tasks.
 Faithfulness—Our theoretical understanding of prompt-based reasoning with LLMs is incomplete.
@@
 In general, more insight into the working of multi-step in-context learning in LLMs is dearly needed.
-Small language models—Efficiency is an important element for wide adoption of language models.
-Important topics are distillation of reasoning to small language models and an understanding of scaling laws.
-Few-shot Reinforcement Learning—Small reasoning problems can be solved with few-shot in-context learning.
-Can we solve larger sequential decision processes, reducing the sample complexity in reinforcement learning?
+Smaller language models—Efficiency is an important element for wide adoption of language models.
+Smaller models have many advantages over larger models: a smaller environmental footprint, more accessible to people and organizations with fewer computational resources, and the possibility to finetune them.
+Models may have fewer parameters, or the parameters may be quantized with fewer bits.
+Unfortunately, with a smaller number of parameters, the models also have less reasoning power.
+It is important for future work to investigate this trade-off.
+Solutions to compensate the limited capabilities of smaller models are distillation of reasoning to small language models and reinforcement learning approaches (Shao et al., 2024) to improve the efficiency of the finetuning/reasoning pipeline (Guo et al., 2025; Du et al., 2025).
 For longer term future work, the following more fundamental questions are important:
-Symbolic and Connectionist Computation—How can we further improve LLM-reasoning: how can LLMs benefit from symbolic reasoning prompts and how can LLMs help ground symbolic reasoning in language?
-Metacognition—Much of the research into reasoning guides the model how it should solve a problem.
-Is it helpful to introduce named concepts for different kinds of reasoning?
-Can the model find these concepts by itself?
-Making the LLM “think” step by step is a first step towards influencing the model’s own “thought” processes.
-The first works on LLM metacognition have appeared, and artificial general intelligence will pursue this further.
-7 Conclusion
+Symbolic and Connectionist Computation—How can we further improve LLM-reasoning: how can LLMs benefit from reasoning prompts based on the symbolic AI literature, and how can LLMs help ground symbolic reasoning in language?
+Multimodal World-Models with Embodied Grounding—How can an LLM maintain a unified, continually updated representation that integrates text, vision, audio, and sensorimotor feedback, enabling it to reason over events across modalities, and refine its world-model through closed perception–action loops in real-world environments?
+Norm-Sensitive and Value-Aligned Reasoning—How do we represent, adapt, and audit diverse cultural norms and ethical values within the reasoning process, ensuring that each step of an LLM’s chain-of-thought respects context-dependent moral constraints while remaining transparent and verifiable?
+Reasoning in other languages than English—The majority of the research into LLM reasoning is for English.
+Although there are efforts in creating benchmarks to test LLM capabilities in other languages555https://github.com/NaiveNeuron/awesome-multilingual-llm-benchmarks, these contain mainly NLP tasks such as question answering (QA) and natural language inference (NLI).
+In addition, challenges related to low-resource languages (languages for which very limited training data is available) have not been addressed yet.
+Reasoning in specialized domains—In an effort to guide development and evaluation of new methods, research in AI has a strong focus on benchmarking: standardized datasets with a limited set of problems that are realistic to evaluate.
+In the real world, however, reasoning problems also occur in more challenging contexts.
+A few examples are: legal reasoning for the interpretation of case law or writing contracts; scientific reasoning for advancing scientific fields; and medical reasoning for AI-assisted diagnostics.
+5. Conclusion
 Prompt-based in-context learning is an efficient machine learning method, requiring no parameter updates to the LLM.
-While achieving good performance on language tasks (System 1), performance on reasoning tasks (System 2) was lacking.
-Reasoning tasks, such as math word problems, are typically solved in step-by-step fashion.
+While achieving good performance on language tasks, performance on reasoning tasks was lacking originally.
+Reasoning tasks, such as math word problems, are typically solved in a step-by-step fashion.
 Recently prompts have been developed that guide an LLM to “think step by step” (Chain-of-thought), and to evaluate and verify the step results.
-The performance of reasoning with LLMs has improved greatly.
+The performance of multi-step reasoning with LLMs has improved greatly, and the field has progressed beyond math word problems.
 Together, the surveyed methods allow the LLM to follow high-quality multi-step reasoning chains.
@@
 For complex reasoning tasks a large number of reasoning steps may be generated.
-To control the size of the reasoning space interactively, external scaffolding algorithms can be used.
-Often, variations on search algorithms or reinforcement learning are used.
+To control the size of the multi-step reasoning space dynamically, external scaffolding algorithms can be used.
+Often, variations on search algorithms are used, and especially reinforcement learning.
 The symbolic and connectionist AI traditions come together in reasoning prompts and search algorithms that help LLM neural networks solve natural language math word and related problems.
+Inference-time reasoning results can be used to augment finetuning, in a feedback loop.
+LLMs hallucinate and suffer from bias, and their use poses different ethical dangers.
+In multi-step reasoning methods, self-verification methods have been developed to reduce error-accumulation, yet ethical dangers may remain, especially since reflection is not error-free.
+Additionally, retrieval augmentation methods ground LLM output directly in sources such as Wikipedia.
 Among the most popular reasoning benchmarks in this survey is GSM8K, which contains 8500 grade school math word problems.
-With LLMs such as GPT-3, reasoning approaches show an improvement of 20-50% points over standard prompting methods.
-For further progress in the field, the development of other challenging benchmarks is important.
-The field of reasoning with LLMs is quite new, and theoretical understanding is lacking in important areas, such as faithful reasoning (models may sometimes find the right answer for the wrong reason).
-Although prompt-based learning allows few-shot learning, the computational needs of LLMs pretraining and finetuning are still high, hence the interest in small language models.
-Reasoning skills that work in large models can often be transferred to small models.
+With older LLMs such as GPT-3, reasoning approaches show an improvement of 20-50% points over standard prompting methods, and some even more.
+Many current LLMs have since adopted the reasoning approaches.
+The success of reasoning with LLMs has attracted more applications, and with them, benchmarks are diverging.
+The field of reasoning with LLMs is quite new, and theoretical understanding is lacking in important areas, such as faithful reasoning (models may sometimes find the right answer using an incorrect reasoning chain).
+There may be more opportunities for efficiency gains by integrating the training of reasoning models, using reinforcement learning, both in training and in context.
+Although prompt-based learning allows few-shot learning at inference time, the computational needs of LLM pretraining and finetuning are still high, hence the interest in small language models.
+Reasoning skills that work in large models can often be distilled to small models.
 Human thought is capable of metacognition, we can think about our thinking process.
@@
 The first preliminary experiments of language models that reason about their reasoning skills have appeared.
-LLM-reasoning is an active field of research, with connections to artificial general intelligence.
-The field has shown great progress.
-Based on current limitations and open questions we provide a research agenda highlighting opportunities for further progress in harder reasoning problems, metacognition, and small language models, amongst others.
+LLM-reasoning is an active field of research that shows great progress.
+Based on current limitations and open questions we provide a research agenda highlighting opportunities for further progress with reinforcement learning (for finetuning, in context reinforcement learning and self-reflection), for LLMs that generate code for external tools, for multimodal world models, for value-aligned reasoning, and for small language models, among others.
+Acknowledgements
+We thank the anonymous reviewers for their valuable suggestions that have considerably improved the paper.